### v5

# Measurement Error and Counterfactuals in Quantitative Trade and Spatial Models

Abstract
Counterfactuals in quantitative trade and spatial models are functions of the current state of the world and the model parameters. Common practice treats the current state of the world as perfectly observed, but there is good reason to believe that it is measured with error. This paper provides tools for quantifying uncertainty about counterfactuals when the current state of the world is measured with error. I recommend an empirical Bayes approach to uncertainty quantification, and show that it is both practical and theoretically justified. I apply the proposed method to the settings in adao2017nonparametric and allen2022welfare and find non-trivial uncertainty about counterfactuals.
1 Introduction
Economists use quantitative trade and spatial models to evaluate counterfactual scenarios. For instance, how do expenditure patterns across countries adjust in response to the implementation of a trade agreement? How are welfare levels affected when transportation infrastructure connecting regions is improved? These counterfactual questions are typically posed relative to an observed factual situation. This implies that the estimand of interest depends directly on the realized data, rather than on the underlying data-generating distribution—a departure from standard statistical settings.
The setting with data-dependent counterfactual estimands is further complicated by the fact that data in quantitative trade and spatial models are often measured with error (goes2023reliability; linsi2023problem; teti2023missing). Unlike classical measurement error settings, where the estimand is typically a parameter of the correctly measured population distribution, here it is a functional of the realized data. To illustrate, consider the canonical Armington model (armington1969theory), where predicted welfare changes from hypothetical trade cost shocks can be written as a function of baseline bilateral trade flows and the trade elasticity (arkolakis2012new). The question I address is how measurement error in the observed trade flows affects uncertainty in the welfare predictions.
I develop an empirical Bayes framework for quantifying uncertainty around counterfactual predictions. The approach requires specifying both a measurement error model and a prior distribution over the latent true data, up to a set of hyperparameters. These hyperparameters are estimated from the observed data via an empirical Bayes step. Bayes’ rule then yields an estimated posterior distribution over the latent data given the noisy observations. Given the structure of quantitative trade and spatial models, this posterior induces a corresponding posterior over counterfactual predictions. Uncertainty can then be summarized by reporting an interval of posterior quantiles.
A natural accompanying point estimator for this procedure is the posterior median, which is always guaranteed to be contained in the reported interval. By contrast, the standard point estimator, which does not account for measurement error, may lay outside the interval. The posterior median answers the question: What does a Bayesian believe the counterfactual prediction would have been in the absence of measurement error? While this is a natural and intuitive question to ask, the answer necessarily depends on the prior.
In settings where the observed data consist of non-negative dyadic flows, I propose a default specification for the measurement error model and prior that can be calibrated directly from the data, yielding a widely applicable empirical Bayes approach. Specifically, I model measurement error as log-normal and use a log-normal prior centered on a structural gravity equation, with a point mass at zero to accommodate zero flows. This setup is designed for ease of implementation and is suitable for a wide class of quantitative trade and spatial models.
I consider two approaches to calibrating the hyperparameters in this default specification. The first assumes constant measurement error variances across flows and relies on researcher input or domain knowledge. The second, applicable in the case of international trade, uses the mirror trade dataset compiled by linsi2023problem, which reports bilateral trade flows as recorded by both exporters and importers. I interpret these paired observations as two independent noisy measurements of the true trade flow, enabling calibration of flow-specific measurement error variances.
To illustrate the impact of incorporating measurement error into counterfactual analysis, I revisit the applications in adao2017nonparametric and allen2022welfare. In adao2017nonparametric, which quantify the welfare effects of China’s accession to the WTO, I model measurement error in baseline bilateral trade flows. I apply the default empirical Bayes approach and construct uncertainty intervals that account for measurement error in the estimated changes in China’s welfare from 1996 to 2011. These intervals are substantially wider than those reported in adao2017nonparametric, which reflect estimation uncertainty.
In the setting of allen2022welfare, the counterfactual question concerns which highway links in the United States yield the highest return on investment and are therefore most promising for improvement. I model measurement error in traffic flows and apply the default empirical Bayes approach, calibrating the prior and measurement error model using estimates from musunuru2019applications. I compute uncertainty intervals that account for measurement error for the three links with the highest estimated returns. Although the intervals are wide, the relative ranking of the top three links remains robust.
This paper contributes to a growing body of work aimed at improving counterfactual analysis in quantitative trade and spatial models (balistreri2008gravity; adao2017nonparametric; kehoe2017quantitative; adao2023putting; ansari2024quantifying; sanders2025new). The most closely related work is DingelTintelnot:2025, which studies calibration procedures in granular environments. That paper considers models that presume a continuum of agents and shows that, when data are limited, unit-level idiosyncrasies are absorbed into the model parameters, leading to overfitting and poor out-of-sample performance. My focus is on the complementary issue of uncertainty quantification due to measurement error—an issue that persists even in non-granular settings. DingelTintelnot:2025 recommends replacing raw observed data with fitted values from a low-dimensional model. I show how this recommendation can be nested into the proposed Bayesian framework.
The remainder of the paper is organized as follows. Section 2 introduces the setting and notation. Section 3 presents the empirical Bayes framework for accounting for measurement error in quantitative trade and spatial models. Section 4 describes a widely applicable default approach. Section 5 demonstrates the procedure in the context of the Armington model. Section 6 applies the method to the trade setting in adao2017nonparametric and explores its use in the economic geography framework of allen2022welfare. Section 7 concludes.
2 Counterfactuals in Quantitative Trade and Spatial Models
This section introduces the notation and discusses the key assumption that commonly underlies counterfactual analyses in quantitative trade and spatial models.
2.1 Notation and Key Assumption
To begin, consider a baseline setting where there is no measurement error. Let $D\in\mathcal{D}\subseteq\mathbb{R}^{d_{D}}$ denote a data vector drawn from distribution $\mathcal{P}_{D}$, and let $\theta\in\Theta\subseteq\mathbb{R}^{d_{\theta}}$ denote a structural parameter. Our objective is to compute a scalar counterfactual quantity $\gamma\in\mathbb{R}$. The key assumption that the counterfactual object of interest has to satisfy is:
Assumption 1.
For a given counterfactual question and fixed parameter value $\theta$, the counterfactual object $\gamma$ can be expressed as a function of the realized data $D$:
$\gamma$ $=g\left(D,\theta\right),$ (1)
for some known function $g:\mathcal{D}\times\Theta\rightarrow\mathbb{R}$.
The exact functional form of $g$ depends on the specific quantitative model that is considered. In Appendix A I discuss Assumption 1 for two leading classes of models, namely invertible models and exact hat algebra models.
The main appeal of focusing on objects of the form in Assumption 1 is that it allows researchers to answer counterfactual questions posed relative to a specific, observed factual situation. In quantitative trade and spatial settings, such questions are often at least as relevant as those concerning average effects. For example, in a quantitative model of international trade, the goal is typically to understand what would happen to the world following a specific policy change, rather than what would occur in a randomly drawn year under that policy.
Assumption 1 implies that if the data $D$ are observed without error and the structural parameter $\theta$ is known, we can perfectly recover $\gamma$. [^1] Indeed, by fixing $g$ I abstract away from model misspecification, an important problem I engage with in future work. This contrasts with standard econometric models, where the object of interest is a function of the correctly measured distribution of the data, rather than the actual observations. So the key distinction with standard settings is:
$\begin{cases}\begin{array}[]{c}\mathrm{standard\ setting:}\\ \mathrm{this\>paper:}\end{array}&\begin{array}[]{c}\gamma=g\left(\mathcal{P}_{D},\theta\right)\\ \gamma=g\left(D,\theta\right),\quad D\sim\mathcal{P}_{D}\end{array}.\end{cases}$ (2)
Importantly, this difference implies that it would not suffice to be able to perfectly estimate the distribution $\mathcal{P}_{D}$. Towards uncertainty quantification, we hence need to account for uncertainty about the realized data themselves rather than their distribution.
3 Empirical Bayes Uncertainty Quantification
This section introduces measurement error into quantitative trade and spatial models. It outlines how to quantify the resulting uncertainty for the counterfactual prediction of interest.
3.1 Prior and Measurement Error Model
Under Assumption 1, our object of interest can be written as a function solely of the data realizations and the structural parameter, which is convenient for answering relevant counterfactual questions. However, the data realizations are economic variables which are often measured with error. For instance, ortiz2018international and goes2023reliability highlight that there are large discrepancies between and within various data sources from trade and international economics. Motivated by this, I assume that, instead of the true data vector $D$, we observe a noisy version $\tilde{D}$.
For uncertainty quantification for the counterfactual prediction, we will require the posterior distribution of the true data given the noisy data. Towards that end, I introduce a model for the measurement error and a prior distribution for the true underlying data,
$\begin{cases}\begin{array}[]{c}\mathrm{prior:}\\ \mathrm{measurement\>error:}\end{array}&\begin{array}[]{c}\pi^{\mathrm{prior}}\left(D|\vartheta\right)\\ \pi^{\mathrm{me}}\left(\tilde{D}|D,\vartheta\right)\end{array}.\end{cases}$
Here, $\vartheta\in\mathbb{R}^{d_{\vartheta}}$ is a vector of unknown hyperparameters.
3.2 Empirical Bayes and Posterior Distribution
Given such a prior distribution and a measurement error model, we can use an empirical Bayes approach to estimate the unknown hyperparameters. [^2] Rather than estimating the parameters of the prior distribution for the true underlying data, which corresponds to an empirical Bayes approach, one could alternatively specify prior distributions for these parameters, which corresponds to a hierarchical Bayes approach. Formally, we have
$\tilde{\vartheta}=\underset{\vartheta}{\arg\max}\int\pi^{\mathrm{me}}\left(\tilde{D}|D,\vartheta\right)\pi^{\mathrm{prior}}\left(D|\vartheta\right)dD.$
Then, given the estimated hyperparameters $\tilde{\vartheta}$, we can use Bayes’ rule to find the estimated posterior distribution of the true data given the noisy data,
$\pi^{\mathrm{post}}\left(D|\tilde{D},\tilde{\vartheta}\right)=\frac{\pi^{\mathrm{me}}\left(\tilde{D}|D,\tilde{\vartheta}\right)\pi^{\mathrm{prior}}\left(D|\tilde{\vartheta}\right)}{\int\pi^{\mathrm{me}}\left(\tilde{D}|D,\tilde{\vartheta}\right)\pi^{\mathrm{prior}}\left(D|\tilde{\vartheta}\right)dD}.$ (3)
Using this estimated posterior, we can generate draws for the true data given the noisy data. [^3] Note that the measurement error distribution does not have to be mean zero, so also allows for measurement error bias. Nevertheless, even mean zero measurement error can result in bias in the counterfactual prediction of interest. This is automatically taken into account by the Bayesian approach when quantifying uncertainty. Furthermore, the individual measurement error distributions can be arbitrarily correlated in this general setup.
The Bayesian approach allows researchers to incorporate economic knowledge through the prior. For example when considering measurement error in non-negative flows between locations, one can fit a prior centered on a gravity model, which I will do in Section 4.
3.3 Quantifying Uncertainty about $\gamma$
[…]
Towards uncertainty quantification, we want to sample from this posterior and report the relevant quantiles. [^4] Note that counterfactual predictions are typically derived as functions of the full system of counterfactual equilibrium variables. Thus, whether the researcher is ultimately interested in a scalar outcome, a relative comparison, or a global average, the mechanics of uncertainty quantification—drawing from the posterior over the true data and solving for equilibrium—remain the same. The entire procedure is summarized in Algorithm 1.
Algorithm 1 Uncertainty quantification about $\gamma=g\left(D,\theta\right)$
  1. 1.
    Input: prior $\pi^{\mathrm{prior}}\left(D|\vartheta\right)$, measurement error model $\pi^{\mathrm{me}}\left(\tilde{D}|D,\vartheta\right)$, noisy data $\tilde{D}$, structural parameter $\theta$, number of bootstrap draws $B$, coverage level $1-\alpha$ (choose $B$ and $\alpha$ such that $\alpha/2\cdot B\in\mathbb{N}$).
  2. 2.
    Empirical Bayes estimation step: $\tilde{\vartheta}=\underset{\vartheta}{\arg\max}\int\pi^{\mathrm{me}}\left(\tilde{D}|D,\vartheta\right)\pi^{\mathrm{prior}}\left(D|\vartheta\right)dD.$
  3. 3.
[…]
  4. 4.
    For b=1,,Bb=1,...,B,
    1. (a)
[…]
    2. (b)
      Compute $\gamma_{b}=g\left(D_{b},\theta\right).$
  5. 5.
    Sort $\left\{\gamma_{b}\right\}_{b=1}^{B}$ to obtain $\left\{\gamma^{(b)}\right\}_{b=1}^{B}$ with $\gamma^{(1)}\leq\gamma^{(2)}\leq...\leq\gamma^{(B)}$.
  6. 6.
    Report $\left[\gamma^{(\alpha/2\cdot B)},\gamma^{(\left(1-\alpha/2\right)\cdot B)}\right].$
A natural accompanying point estimator for the procedure in Algorithm 1 is the posterior median, which is always guaranteed to be contained in the reported interval. By contrast, the standard point estimator $g\left(\tilde{D},\theta\right)$, which does not account for measurement error, may lay outside the interval. The posterior median corresponds to an optimal estimate under the estimated posterior and under absolute value loss from a decision-theoretic perspective (see for example Proposition 2.5.5 in robert2007bayesian).
Note that the posterior median does not guarantee bias correction in the frequentist sense. It answers the question: What does a Bayesian believe the counterfactual prediction would have been in the absence of measurement error? While this is a natural and intuitive question to ask, the answer necessarily depends on the prior. But when the prior reflects well-established economic relationships—such as gravity patterns in quantitative trade and spatial models—the posterior median provides a principled estimate of the counterfactual prediction.
3.4 Relation to the Literature
3.4.1 Relation to Measurement Error Literature
The literature on measurement error in nonlinear models is extensive, as reviewed in hu2015microeconomic and schennach2016recent, and the most closely related strand of measurement error literature is that on nonseparable error models (matzkin2003nonparametric; chesher2003identification; hoderlein2007identification; matzkin2008identification; hu2008instrumental; schennach2012local; song2015estimating). However, these results do not apply to my setting.
The key distinguishing feature of the setting in this paper is that the object of interest $\gamma$ directly depends on the correctly measured data, because the equality in Assumption 1 is an exact statement. This is convenient for answering counterfactual questions, and arises because counterfactual questions are typically posed relative to an observed factual situation. In contrast, in standard econometric methods of measurement error, the object of interest is a function of the correctly measured distribution of the data, $\mathcal{P}_{D}$, rather than the actual realized observations, $D$. This leads to the key distinction in Equation (2).
This difference is important because in my setting, it would not suffice to be able to perfectly estimate the distribution $\mathcal{P}_{D}$. For example in a quantitative model of international trade, to answer counterfactual questions we need the realized trade flows, rather than the trade flow distribution from which they are drawn. In contrast, in standard econometric models of measurement error, knowing this distribution would suffice, because the estimands are functionals of the correctly measured distribution of the data. By virtue of that, we need to account for uncertainty about the observations themselves rather than their distribution.
3.4.2 Relation to DingelTintelnot:2025
The most relevant paper in the literature on improving counterfactual calculations in quantitative trade and spatial economics is DingelTintelnot:2025, which studies calibration procedures in granular settings. In these settings, individual idiosyncrasies do not wash out and can cause overfitting and poor performance out-of-sample. To deal with this, DingelTintelnot:2025 proposes to, instead of the observed data, either use fitted values obtained using a low-dimensional model or smooth the data using matrix approximation techniques. Both of these approaches can be cast as special cases of the proposed procedure in Algorithm 1, by choosing a specific prior.
Specifically, the main recommendation is to use a low-dimensional model and is called the covariates-based approach. DingelTintelnot:2025 considers a quantitative spatial model with $L$ individuals. Let $\ell_{ij}$ denote the share of people residing in location $i$ and working in location $j$. The covariates-based approach then interprets the observed commuting shares $\left\{\tilde{\ell}_{ij}\right\}$ as a finite sample from a continuum model. This results in the maximum likelihood model
$\left\{\tilde{\ell}_{ij}\cdot L\right\}|\vartheta\sim\mathrm{Multinomial}\left(\left\{h_{ij}\left(\vartheta\right)\right\}\right),$ (4)
for $\vartheta$ a set of hyperparameters and $h_{ij}\left(\vartheta\right)$ a model function which I discuss further in Appendix B.
[…]
$\tilde{\gamma}^{DT}=g\left(\left\{h_{ij}\left(\tilde{\vartheta}\right)\right\},\theta\right)$
instead of $g\left(\left\{\tilde{\ell}_{ij}\right\},\theta\right)$.
To see how the covariates-based approach from DingelTintelnot:2025 is nested in my Bayesian framework, consider the following prior and measurement error model,
$\begin{cases}\begin{array}[]{c}\mathrm{prior:}\\ \mathrm{measurement\>error:}\end{array}&\begin{array}[]{c}\ell_{ij}|\vartheta\sim\delta_{h_{ij}\left(\vartheta\right)},\quad i,j=1,...,n\\ \left\{\tilde{\ell}_{ij}\cdot L\right\}|\vartheta\sim\mathrm{Multinomial}\left(\left\{\ell_{ij}\right\}\right)\end{array},\end{cases}$ (5)
where $\delta_{h_{ij}\left(\vartheta\right)}$ denotes the Dirac mass at $h_{ij}\left(\vartheta\right)$, implying a degenerate prior. The empirical Bayes step then combines the prior and measurement error model to find
$\left\{\tilde{\ell}_{ij}\cdot L\right\}|\vartheta\sim\mathrm{Multinomial}\left(\left\{h_{ij}\left(\vartheta\right)\right\}\right),$
which overlaps with the model in Equation (4), and uses maximum likelihood estimation to estimate $\vartheta$ by $\tilde{\vartheta}$. This yields the estimated prior and measurement error model
$\begin{cases}\begin{array}[]{c}\mathrm{prior:}\\ \mathrm{measurement\>error:}\end{array}&\begin{array}[]{c}\ell_{ij}|\tilde{\vartheta}\sim\delta_{h_{ij}\left(\tilde{\vartheta}\right)},\quad i,j=1,...,n\\ \left\{\tilde{\ell}_{ij}\cdot L\right\}|\tilde{\vartheta}\sim\mathrm{Multinomial}\left(\left\{\ell_{ij}\right\}\right)\end{array}.\end{cases}$
Using Bayes’ rule we can then find the estimated posterior for the true commuting shares,
$\ell_{ij}|\left\{\tilde{\ell}_{ij}\right\},\tilde{\vartheta}\sim\delta_{h_{ij}\left(\tilde{\vartheta}\right)},\quad i,j=1,...,n.$ (6)
Note that after the empirical Bayes estimation step, since the estimated prior is degenerate, no information is taken from the estimated measurement error model.
The estimated posterior distributions in Equation (6) translate to an estimated posterior for $\gamma$,
$\pi^{\mathrm{post}}\left(\gamma|\left\{\tilde{\ell}_{ij}\right\},\tilde{\vartheta}\right)=\delta_{g\left(\left\{h_{ij}\left(\tilde{\vartheta}\right)\right\},\theta\right)}.$
[…]
Note that if we follow Algorithm 1 exactly, then in step 4a—where we draw from the posterior distributions in Equation (6)—we will obtain the same values in each bootstrap iteration. As a result, the interval constructed in step 6 will collapse to a single point. This outcome is unsurprising, as the procedure in DingelTintelnot:2025 focuses on point estimation and correcting overfitting bias, rather than on quantifying estimation uncertainty. Accordingly, if the sole concern is overfitting bias, their method provides an appropriate approach.
The second recommendation in DingelTintelnot:2025 is to replace the observed data with a smoothed version using matrix approximation techniques. I discuss this approach in Appendix B.
3.5 Estimation Error
The counterfactual prediction of interest will typically depend on a structural parameter $\theta$. It is common in applied work to plug in a fixed value for the structural parameter taken from the literature or obtained through data-driven methods, thus ignoring the uncertainty associated with the estimation process. An exception is adao2017nonparametric, which reports confidence sets for the counterfactual predictions of interest that account for estimation error.
Towards accounting for estimation error for quantitative trade and spatial models in the presence of measurement error, let $\tilde{\theta}$ denote the estimator of the estimand $\theta$. This estimand is usually a function of the distribution of the data $\mathcal{P}_{D}$. This implies that, to address measurement error affecting the structural parameter, one can apply the frequentist measurement error techniques discussed in Section 3.4.1 to find a bias-corrected estimate, though the resulting correction will not admit a Bayesian interpretation.
Alternatively, in Appendix C I outline a fully Bayesian approach that also considers estimation error. Specifically, I assume that the posterior distribution of the structural parameter $\theta$ given the true data $D$ is approximately normal, which is justified under regularity conditions that are closely related to those required for frequentist asymptotic normality. We then have two different posteriors,
[…]
As in Section 3.3, a natural point estimator for the structural parameter is the median of the estimated posterior given the noisy data,
$\pi^{\mathrm{post}}\left(\theta|\tilde{D},\tilde{\vartheta}\right)=\int\pi^{\mathrm{post,ee}}\left(\theta|D\right)\pi^{\mathrm{post,me}}\left(D|\tilde{D},\tilde{\vartheta}\right)dD.$
Using Assumption 1, we can also find the posterior $\pi^{\mathrm{post,ee}}\left(\gamma|D\right)$, and it follows that a natural point estimator for the counterfactual prediction is the median of the estimated posterior
[…]
However, for models satisfying Assumption 1, I am not aware of a frequentist framework that jointly incorporates estimation error and measurement error within a unified procedure that permits both point estimation and uncertainty quantification. By contrast, the proposed Bayesian approach accommodates both sources of uncertainty within a single coherent framework. Moreover, it nests the case in which only measurement error is present: as estimation error vanishes, the procedure naturally reduces to the framework that accounts solely for measurement error. I therefore recommend the Bayesian approach when both sources of uncertainty are relevant, and either the Bayesian or frequentist approach when only estimation error is of concern. In the latter case, the proposed framework can still serve as a robustness check with respect to measurement error.
4 Widely Applicable Default Approach
This section proposes a default empirical Bayes approach that can be applied in many settings, as it is both economically reasonable for many quantitative trade and spatial models and computationally convenient. It also discusses the toolkit that accompanies the paper.
4.1 Default Prior and Measurement Error Model
In many applications, the support of the true data and the sampling process naturally suggest both a measurement error model and a prior. For example, when observing migration shares constructed from count data, a multinomial measurement error model arises naturally from sampling variation. Since the true shares are non-negative and must sum to one, a Dirichlet prior provides a coherent and tractable specification.
For settings where such natural choices are not available, this section proposes a widely applicable default approach for quantifying uncertainty about the counterfactual prediction of interest. This default approach can be applied out-of-the-box to many quantitative trade and spatial models, but can also be easily adapted to other settings. It recommends default choices for the prior distribution and measurement error model, and discusses how to calibrate both based on observed data.
Concretely, consider the setting where we can write $\gamma=g\left(\left\{F_{ij}\right\},\theta\right)$, for $\left\{F_{ij}\right\}$ a set of non-negative flows between locations. This setup is commonplace in quantitative trade and spatial models (costinot2014trade; redding2017quantitative; proost2019can). I assume that both the prior distributions on the true flows and the measurement errors are mixtures of a point mass at zero and a log-normal distribution, a so-called spike-and-slab distribution (mitchell1988bayesian). The point mass at zero is necessary because in both trade and spatial applications bilateral flows of zeros are common, particularly when considering more granular data (helpman2008estimating; DingelTintelnot:2025). This prior and measurement error model imply that the posterior distribution of the true flows given the noisy flows will also be a spike-and-slab distribution. This mixture model is fairly flexible, and conjugacy ensures computational tractability. Furthermore, I assume that the prior mean exhibits a gravity relationship, for which there is strong empirical evidence (head2014gravity; allen201813). [^5] One can easily enrich this gravity prior by adding other “distance” variables such as differences in income or productivity, or by adding dummies that indicate similarity such as contiguity or a common language, see for example silva2006log. I experimented with this but the results do not change much. This is summarized in the following assumption:
Assumption 2.
We have
$\begin{cases}\begin{array}[]{c}\mathrm{true\ zeros:}\\ \mathrm{spurious\ zeros:}\\ \mathrm{prior:}\\ \\ \mathrm{likelihood:}\\ \\ \end{array}&\begin{array}[]{c}P_{ij}\sim\mathrm{Bern}\left(p_{ij}\right)\\ B_{ij}\sim\mathrm{Bern}\left(b_{ij}\right)\\ F_{ij}\sim P_{ij}\cdot\delta_{0}+\left(1-P_{ij}\right)\cdot e^{\mathcal{N}\left(\mu_{ij},s_{ij}^{2}\right)}\\ \mu_{ij}=\beta\log\mathrm{dist}_{ij}+\alpha_{i}^{\mathrm{orig}}+\alpha_{j}^{\mathrm{dest}}\\ \tilde{F}_{ij}|F_{ij}\sim\delta_{0}\cdot\mathbb{I}\left\{F_{ij}=0\right\}\\ +\left[B_{ij}\cdot\delta_{0}+\left(1-B_{ij}\right)\cdot e^{\mathcal{N}\left(\log F_{ij},\varsigma_{ij}^{2}\right)}\right]\cdot\mathbb{I}\left\{F_{ij}>0\right\}\end{array},\end{cases}$
for i,j=1,,ni,j=1,...,n, where $\delta_{0}$ denotes the Dirac mass at zero, $\mathrm{dist}_{ij}$ denotes the distance between locations $i$ and $j$, $\alpha_{i}^{\mathrm{orig}}$ is an origin fixed effect and $\alpha_{j}^{\mathrm{dest}}$ is a destination fixed effect.
The probability that a bilateral trade flow is truly zero is denoted by $p_{ij}$, and a true zero flow is assumed to always result in an observed zero. [^6] Assumption 2 implies that both true and spurious zeros occur randomly. Alternatively, one could think about incorporating endogenous zeros using selection mechanisms such as in helpman2008estimating. The probability of a spurious zero—that is, an observed zero despite a non-zero underlying true flow—is denoted by $b_{ij}$. The prior means and variances are denoted by $\left\{\mu_{ij}\right\}$ and $\left\{s_{ij}^{2}\right\}$, respectively. The flow-specific measurement error variances are denoted by $\left\{\varsigma_{ij}^{2}\right\}$.
Gather the hyperparameters in $\vartheta=\left(\left\{p_{ij}\right\},\left\{b_{ij}\right\},\beta,\left\{\alpha_{i}^{\mathrm{orig}}\right\},\left\{\alpha_{i}^{\mathrm{dest}}\right\},\left\{s_{ij}^{2}\right\},\left\{\varsigma_{ij}^{2}\right\}\right)$. It follows that the posterior distribution for the true flow between location $i$ and $j$, $F_{ij}$, given its noisy version, $\tilde{F}_{ij}$ is given by
$F_{ij}|\tilde{F}_{ij},\vartheta\sim\begin{cases}\begin{array}[]{c}Q_{ij}\cdot\delta_{0}+\left(1-Q_{ij}\right)\cdot e^{\mathcal{N}\left(\mu_{ij},s_{ij}^{2}\right)}\\ \exp\left\{\mathcal{N}\left(\frac{s_{ij}^{2}}{s_{ij}^{2}+\varsigma_{ij}^{2}}\log\tilde{F}_{ij}+\frac{\varsigma_{ij}^{2}}{s_{ij}^{2}+\varsigma_{ij}^{2}}\mu_{ij},\left(\frac{1}{s_{ij}^{2}}+\frac{1}{\varsigma_{ij}^{2}}\right)^{-1}\right)\right\}\end{array}&\begin{array}[]{c}\tilde{F}_{ij}=0\\ \tilde{F}_{ij}>0\end{array},\end{cases}$ (7)
for i,j=1,,ni,j=1,...,n, where $Q_{ij}\sim\mathrm{Bern}\left(\frac{p_{ij}}{p_{ij}+b_{ij}\left(1-p_{ij}\right)}\right)$. [^7] Note that $F_{ij}$ are drawn independently across $ij$-pairs. As a result, certain adding-up constraints may no longer hold. In principle, this issue could be addressed by drawing from a constrained joint distribution.
Conditional on being able to calibrate the parameters $\vartheta$—the empirical Bayes estimation step—one can quantify uncertainty about $\gamma$ by finding the interval as described in Algorithm 1. Then, a default procedure for quantifying uncertainty about $\gamma$ is summarized in Algorithm 2.
Algorithm 2 Uncertainty quantification about $\gamma=g\left(\left\{F_{ij}\right\},\theta\right)$
  1. 1.
    Input: noisy flows $\left\{\tilde{F}_{ij}\right\}$, structural parameter $\theta$, number of bootstrap draws $B$, coverage level $1-\alpha$ (choose $B$ and $\alpha$ such that $\alpha/2\cdot B\in\mathbb{N}$).
  2. 2.
    Empirical Bayes estimation step: calibrate $\vartheta$ as outlined in Section 4.2 and denote the estimator by $\tilde{\vartheta}$.
  3. 3.
    For b=1,,Bb=1,...,B,
    1. (a)
[…]
    2. (b)
      Compute $\gamma_{b}=g\left(\left\{F_{ij,b}\right\}_{i,j=1}^{n},\theta\right)$.
  4. 4.
    Sort $\left\{\gamma_{b}\right\}_{b=1}^{B}$ to obtain $\left\{\gamma^{(b)}\right\}_{b=1}^{B}$ with $\gamma^{(1)}\leq\gamma^{(2)}\leq...\leq\gamma^{(B)}$.
  5. 5.
    Report $\left[\gamma^{(\alpha/2\cdot B)},\gamma^{(\left(1-\alpha/2\right)\cdot B)}\right].$
Remark 1.
One can verify how reasonable the normality assumption on the prior and measurement error model is by comparing the histogram of the normalized residuals
$\left\{\frac{\log\tilde{F}_{ij}-\left\{\tilde{\beta}\log\mathrm{dist}_{ij}+\tilde{\alpha}_{i}^{\mathrm{orig}}+\tilde{\alpha}_{j}^{\mathrm{dest}}\right\}}{\sqrt{\tilde{s}_{ij}^{2}+\tilde{\varsigma}_{ij}^{2}}}\right\}$
with the probability density function of a standard normal distribution. To further check the reasonableness of the gravity prior, we can look at the adjusted R-squared of the gravity model and, following allen201813, plot the log flows against the log distance for positive flows, after partitioning out the origin and destination fixed effects. In Appendices D and E I perform both these checks for my applications.
Remark 2.
One might be worried about misspecification of the prior and measurement error model. For the normal-normal model, we can use prior density-ratio classes to find worst-case bounds on posterior quantiles over a neighborhood that contains distributions that are not too far away from the assumed normal distribution for the prior and measurement error model. It turns out that incorporating uncertainty around the prior and measurement error model amounts to reporting slightly wider quantiles. The details can be found in Appendix F.
4.2 Empirical Bayes Estimation Step: Calibrating $\vartheta$
The hyperparameters in $\vartheta$ need to be calibrated. I consider two cases.
4.2.1 Baseline Case with Domain Knowledge
In the baseline case I restrict the measurement error variances and prior variances to be constant across flows so that $\varsigma_{ij}^{2}=\varsigma^{2}$ and $s_{ij}^{2}=s^{2}$ for all i,j=1,,ni,j=1,...,n. Furthermore, I require knowledge of the common measurement error variance $\varsigma^{2}$ and of the Bernoulli parameters $\left\{p_{ij}\right\}$ and $\left\{b_{ij}\right\}$. [^8] In the absence of a prior on the measurement error variance, one could adopt a sensitivity analysis approach by plotting intervals while varying the variance. Alternatively, one could determine the minimum level of measurement error that would overturn the counterfactual conclusion. It then remains to estimate $\left(\beta,\left\{\alpha_{i}^{\mathrm{orig}}\right\},\left\{\alpha_{i}^{\mathrm{dest}}\right\},s^{2}\right)$. Towards this, we can combine the equations in Assumption 2 to find
$\log\tilde{F}_{ij}\sim\mathcal{N}\left(\beta\log\mathrm{dist}_{ij}+\alpha_{i}^{\mathrm{orig}}+\alpha_{j}^{\mathrm{dest}},s^{2}+\varsigma^{2}\right),\quad\tilde{F}_{ij}>0.$
Using maximum likelihood estimation, it follows that the prior mean parameters can be estimated from the regression
$\log\tilde{F}_{ij}=\beta\log\mathrm{dist}_{ij}+\alpha_{i}^{\mathrm{orig}}+\alpha_{j}^{\mathrm{dest}}+\phi_{ij},\quad\tilde{F}_{ij}>0,$
with $\phi_{ij}$ an error term. It follows that the estimated prior means and variance are
$\tilde{\mu}_{ij}$ $=\tilde{\beta}\log\mathrm{dist}_{ij}+\tilde{\alpha}_{i}^{\mathrm{orig}}+\tilde{\alpha}_{j}^{\mathrm{dest}},\quad i,j=1,...,n$ (8)
$\tilde{s}^{2}$ $=\max\left\{\widetilde{\operatorname{Var}}\left(\log\tilde{F}_{ij}-\tilde{\mu}_{ij}|\tilde{F}_{ij}>0\right)-\tilde{\varsigma}^{2},0\right\}.$ (9)
Obtaining estimators for these prior means and variances is what walters2024empirical calls the deconvolution step.
4.2.2 Mirror Trade Data
When the non-negative bilateral flows correspond to trade flows between countries, I use the mirror trade dataset from linsi2023problem to calibrate $\vartheta$. This dataset has two estimates of each bilateral trade flow, both as reported by the exporter and as by the importer. linsi2023problem shows that there are so-called mirror discrepancies in bilateral trade flows between almost all countries. This means that, for instance, while the value that Germany reports it imported from France and the value that France reports it exported to Germany should be the same, in practice they are often different. I interpret this as observing two independent noisy observations per time period for each bilateral trade flow. The key identifying assumptions are that the flow-specific probabilities of true zeros, the flow-specific probabilities of spurious zeros, and the flow-specific measurement error variances are constant over time.
The details for the calibration can be found in Appendix G. I first calibrate the probabilities of true zeros $\left\{p_{ij}\right\}$ and the probabilities of spurious zeros $\left\{b_{ij}\right\}$ by noting that for each bilateral trade flow we can use the time variation to identify the probabilities of observing a certain number of zeros. I then leverage the model structure to calibrate the measurement error variances $\left\{\varsigma_{ij}^{2}\right\}$. Lastly, I calibrate the prior parameters, using a similar approach as for the baseline case with domain knowledge. To leverage country information and the fact that importers and exporters can differ in their reliability, I shrink the measurement error and prior variances using country-origin and country-destination fixed effects.
4.3 Toolkit
Accompanying the paper, I provide an easy-to-use toolkit that consists of three programs. [^9] The toolkit is written in MATLAB and can be found on my website, https://sandersbas.github.io/. A version in R is available upon request. The first program implements the high-level approach in Algorithm 1. It takes as inputs $\left(B,\theta,\tilde{D},\pi^{\mathrm{post}}\left(D|\tilde{D},\tilde{\vartheta}\right),g\right)$ and outputs posterior draws $\left\{\gamma_{b}\right\}_{b=1}^{B}$. The second program implements the default approach in Algorithm 2. It takes as inputs $\left(B,\theta,\left\{\tilde{F}_{ij}\right\},\tilde{\vartheta},g\right)$ and again outputs posterior draws $\left\{\gamma_{b}\right\}_{b=1}^{B}$. The third program, which can serve as an input to the second, uses the mirror trade dataset of linsi2023problem and allows the researcher to choose countries and years for which they want to estimate the hyperparameters of the prior and measurement error model. This is summarized in Algorithm 3.
Algorithm 3 Toolkit
  1. 1.
    Program 1: General algorithm.
    • Input: number of draws $B$, structural parameter $\theta$, data $\tilde{D}$, functions $\tilde{D}\mapsto D_{b}$, $\left(D,\theta\right)\mapsto\gamma$.
    • Output: posterior draws $\left\{\gamma_{b}\right\}_{b=1}^{B}$.
  2. 2.
    Program 2: Default approach.
    • Input: number of draws $B$, structural parameter $\theta$, noisy flows $\left\{\tilde{F}_{ij}\right\}$, estimated hyperparameters $\tilde{\vartheta}$, function $\left(\left\{F_{ij}\right\},\theta\right)\mapsto\gamma$.
    • Output: posterior draws $\left\{\gamma_{b}\right\}_{b=1}^{B}$, plot that compares histogram of the normalized residuals with the probability density function of a standard normal distribution as per Remark 1.
  3. 3.
    Program 3: Mirror trade data calibration.
    • Input: countries $\mathcal{I}$, years to produce bootstrap draws for $\mathcal{T}$, years to use for calibration $\mathcal{T}_{\mathrm{calibration}}$.
    • Output (can serve as input to Program 2): noisy flows $\left\{\tilde{F}_{ij}\right\}$, estimated hyperparameters $\tilde{\vartheta}$, adjusted R-squared of the gravity model for the last year in $\mathcal{T}$, plot of log flows against log distance for positive flows, after partitioning out the origin and destination fixed effects as per Remark 1.
5 Prototypical Example: Armington Model
This section illustrates the proposed procedure using the Armington model (armington1969theory), a canonical workhorse model in international trade, as outlined, for example, in costinot2014trade.
5.1 Model and Counterfactual Question of Interest
Countries are indexed by i,j=1,,ni,j=1,...,n, and with CES preferences and perfect competition, it follows that the relevant gravity equations and budget constraints are:
$F_{ij}$ $=\frac{\left(\tau_{ij}Y_{i}\right)^{-\varepsilon}\chi_{ij}}{\sum_{k}\left(\tau_{kj}Y_{k}\right)^{-\varepsilon}\chi_{kj}}E_{j},\qquad i,j=1,...,n$ (10)
$E_{i}$ $=\left(1+\kappa_{i}\right)Y_{i},\qquad i=1,...,n.$ (11)
Here, $F_{ij}$ denotes the trade flow from country $i$ to $j$, and $Y_{i}=\sum_{\ell=1}^{n}F_{i\ell}$, $E_{i}=\sum_{k=1}^{n}F_{ki}$ and $\kappa_{i}=\left(E_{i}-Y_{i}\right)/Y_{i}$ denote country $i$’s total income, total expenditure and the ratio of the trade deficit to income, respectively. Furthermore, $\tau_{ij}$ denotes the iceberg trade cost between country $i$ and $j$, which means that in order to sell one unit of a good in country $j$, country $i$ must ship $\tau_{ij}\geq 1$ units, with $\tau_{ii}=1$. Lastly, $\varepsilon>0$ is the trade elasticity and $\left\{\chi_{ij}\right\}$ are idiosyncratic preferences.
Now, say we are interested in the counterfactual where we change the trade costs $\left\{\tau_{ij}\right\}$ proportionally by $\left\{\tau_{ij}^{\mathrm{cf,prop}}\right\}$, holding the trade elasticity $\varepsilon$, the idiosyncratic preferences $\left\{\chi_{ij}\right\}$ and the trade imbalance variables $\left\{\kappa_{i}\right\}$ constant. In Appendix H.1 I show that we can then solve for the corresponding proportional changes in income, $\left\{Y_{i}^{\mathrm{cf,prop}}\right\}$, using
$Y_{i}^{\mathrm{cf,prop}}Y_{i}=\sum_{j}\frac{\left(\tau_{ij}^{\mathrm{cf,prop}}Y_{i}^{\mathrm{cf,prop}}\right)^{-\varepsilon}}{\sum_{k}\lambda_{kj}\left(\tau_{kj}^{\mathrm{cf,prop}}Y_{k}^{\mathrm{cf,prop}}\right)^{-\varepsilon}}\lambda_{ij}\left(1+\kappa_{j}\right)Y_{j}^{\mathrm{cf,prop}}Y_{j},\qquad i=1,...,n,$
where $\lambda_{ij}=F_{ij}/E_{j}$ denotes the expenditure share that country $j$ spends on goods from country $i$. By Walras’ Law, the proportional changes in income are only pinned down up to a multiplicative constant. Subsequently, following costinot2014trade, we can exactly solve for proportional changes in expenditure shares and welfare (real consumption) levels:
$\lambda_{ij}^{\mathrm{cf,prop}}$ $=\frac{\left(\tau_{ij}^{\mathrm{cf,prop}}Y_{i}^{\mathrm{cf,prop}}\right)^{-\varepsilon}}{\sum_{k}\lambda_{kj}\left(\tau_{kj}^{\mathrm{cf,prop}}Y_{k}^{\mathrm{cf,prop}}\right)^{-\varepsilon}},\qquad i,j=1,...,n$
$W_{i}^{\mathrm{cf,prop}}$ $=\left(\lambda_{ii}^{\mathrm{cf,prop}}\right)^{-1/\varepsilon},\qquad i=1,...,n.$
The income levels $\left\{Y_{i}\right\}$, the expenditure shares $\left\{\lambda_{ij}\right\}$ and the trade deficit variables $\left\{\kappa_{i}\right\}$ are all functions of the trade flows $\left\{F_{ij}\right\}$, so the relevant counterfactual mapping is
$\left\{F_{ij}\right\},\left\{\tau_{ij}^{\mathrm{cf,prop}}\right\},\varepsilon\mapsto\left\{W_{i}^{\mathrm{cf,prop}}\right\}.$
It follows that for a given counterfactual question as described by $\left\{\tau_{ij}^{\mathrm{cf,prop}}\right\}$, we only require knowledge of the baseline trade flows $\left\{F_{ij}\right\}$ and the trade elasticity $\varepsilon$. So we have:
$D$ $=\left\{F_{ij}\right\}$
$\theta$ $=\varepsilon.$
The specific counterfactual question I consider is a 10% increase in all bilateral trade costs between 76 countries, so that $\tau_{ij}^{\mathrm{cf,prop}}=1+0.1\cdot\mathbb{I}\left\{i\neq j\right\}$ for i,j=1,,ni,j=1,...,n. I focus on the proportional percentage changes in welfare in the Central African Republic, the Netherlands, Sweden and the United States. It follows that, fixing $\left\{\tau_{ij}^{\mathrm{cf,prop}}\right\}$, we have
$\gamma_{q}=100\cdot\left(W_{q}^{\mathrm{cf,prop}}-1\right)\equiv g_{q}\left(\left\{F_{ij}\right\},\varepsilon\right),$ (12)
for each $q\in\left\{\mathrm{CAF},\mathrm{NLD},\mathrm{SWE},\mathrm{USA}\right\}$.
5.2 Measurement Error Model and Prior
For the Armington model, I will consider measurement error in trade flows $\left\{F_{ij}\right\}$. Hence, instead of the true trade flows we observe noisy trade flows $\left\{\tilde{F}_{ij}\right\}$, which in turn lead to noisy counterfactual predictions $\tilde{\gamma}_{q}$ for $q\in\left\{\mathrm{CAF},\mathrm{NLD},\mathrm{SWE},\mathrm{USA}\right\}$. Our goal is to quantify the uncertainty that arises from measurement error, and report an accompanying point estimator that aims to answer the question what a Bayesian believes the counterfactual predictions would have been in the absence of measurement error.
[…]
5.3 Results
We can see the impact of measurement error in Table 1 and Figure 1. In Table 1 I compare the standard point estimates based on noisy flows and the posterior median estimates, and report the intervals obtained using Algorithm 2. In Figure 1 I plot the standard point estimates and the smoothed estimated posterior distributions.
We observe that for the Central African Republic and the Netherlands there is a considerable difference between the point estimate and the median of the posterior distribution, causing the point estimate to lie outside the credible set. For Sweden and the United States there is less of a discrepancy. These plots illustrate that the proposed approach automatically incorporates bias induced by measurement error, and that this bias can be both negative and positive. In Appendix H.2, I show the results for all 76 countries in the sample.
[…]
$\gamma_{\mathrm{CAF}}$ -1.09 -0.26 [-0.42, -0.14]
$\gamma_{\mathrm{NLD}}$ -5.15 -6.57 [-7.10, -6.04]
$\gamma_{\mathrm{SWE}}$ -3.25 -3.51 [-3.79, -3.25]
$\gamma_{\mathrm{USA}}$ -1.07 -1.01 [-1.27, -0.56]
[…]
6 Applications
In this section I discuss the applications in adao2017nonparametric and allen2022welfare. The application in adao2017nonparametric extends the prototypical example to a panel-data setting and recovers changes in trade costs using variation over time. The application in allen2022welfare illustrates how the proposed procedure can be extended to an economic geography framework.
6.1 Application 1: adao2017nonparametric
6.1.1 Model and Counterfactual Question of Interest
The empirical application of adao2017nonparametric investigates the effects of China joining the WTO, the so-called China shock. Specifically, the authors examine what would have happened to China’s welfare if China’s trade costs had stayed constant at their 1995 levels. They consider $n$ countries and $T$ time periods. The exercise I am considering is assessing the sensitivity of counterfactual predictions to measurement error in bilateral trade flows.
The counterfactual objects of interest is the change in China’s welfare, defined as the percentage change in income that the representative agent in China would be indifferent about accepting instead of the counterfactual change where China’s trade costs are fixed at their 1995 levels. The details of the model can be found in Appendix D.1.101010In adao2017nonparametric, the authors consider two demand systems: standard CES and “Mixed CES”. I focus on the standard CES specification. The key insight is that we can express the proportional change in China’s welfare in period $t$, denoted $W_{\mathrm{China},t}^{\mathrm{cf,prop}}$, as a function of the full set of bilateral trade flows across periods, $\left\{F_{ij,t}\right\}$, and the trade elasticity, $\varepsilon$. Implementing this mapping requires multiple years of data, since answering the counterfactual question necessitates recovering the proportional change in China’s trade costs between 1995 and year $t$. This change is inferred from bilateral trade flows observed in those two years. Hence, we can write
$W_{\mathrm{China},t}^{\mathrm{cf,prop}}=g_{t}\left(\left\{F_{ij,t}\right\},\varepsilon\right),$ (13)
for t=1,,Tt=1,...,T and known functions $g_{t}:\mathbb{R}_{+}^{Tn\left(n-1\right)}\times\mathbb{R}_{++}\rightarrow\mathbb{R}$. Then, conditional on a prior distribution for the true bilateral flows $\left\{F_{ij,t}\right\}$ and a measurement error model, we can quantify uncertainty for $\left\{W_{\mathrm{China},t}^{\mathrm{cf,prop}}\right\}$.
6.1.2 Measurement Error Model and Prior
The default approach from Section 4 can be applied. For the empirical Bayes step, the calibration of $\vartheta$, we can use the mirror trade data setting from Section 4.2.2. Since there are no zero flows in this application, the estimated posterior of interest is
$F_{ij,t}|\tilde{F}_{ij,t}\sim\exp\left\{\mathcal{N}\left(\frac{\mathring{s}_{ij}^{2}}{\mathring{s}_{ij}^{2}+\mathring{\varsigma}_{ij}^{2}}\log\left(\tilde{F}_{ij,t}\right)+\frac{\mathring{\varsigma}_{ij}^{2}}{\mathring{s}_{ij}^{2}+\mathring{\varsigma}_{ij}^{2}}\tilde{\mu}_{ij,t},\left(\frac{1}{\mathring{s}_{ij}^{2}}+\frac{1}{\mathring{\varsigma}_{ij}^{2}}\right)^{-1}\right)\right\},$
where $\left\{\mathring{s}_{ij}^{2}\right\}$, $\left\{\mathring{\varsigma}_{ij}^{2}\right\}$, $\left\{\tilde{F}_{ij,t}\right\}$ and $\left\{\tilde{\mu}_{ij,t}\right\}$ are all defined in Appendix G.
6.1.3 Results
Having obtained a posterior distribution for the true trade flows given the noisy trade flows, we can now quantify uncertainty about the counterfactual predictions of interest. In Figure 2, I reproduce Figure 3 of adao2017nonparametric, which plots the percentage change in China’s welfare as a result of the China shock for each year in the period 1996-2011, and include two $95\%$ intervals.
The first region only considers estimation error and hence assumes the data are perfectly measured. It is constructed using code provided by the authors, and samples from the normal distribution with mean and variance equal to the GMM estimator for the trade elasticity $\varepsilon$ and its sampling variance, respectively. The resulting intervals are small for the period before the year 2000, and then slowly become wider. These are the intervals reported in adao2017nonparametric.
The second region considers only measurement error and no estimation error in $\varepsilon$. The resulting intervals are considerably wider than the intervals based on estimation error, especially in the first few years. In Appendix D.3 I provide additional discussion and analyses.
[…]
6.2 Application 2: allen2022welfare
6.2.1 Model and Counterfactual Question of Interest
The empirical application in allen2022welfare aims to estimate the returns on investment for all highway segments of the US Interstate Highway network. The authors do so by introducing an economic geography model and calculating what happens to welfare after a $1\%$ improvement to all highway links. Combining these counterfactual welfare changes with how many lane-miles must be added in order to achieve the $1\%$ improvement, they find the highway segments with the greatest return on investment.
This exercise only requires data on incomes and traffic flows of the $n$ locations and knowledge of four structural model parameters. The details of the model can be found in Appendix E.1, but the key relation is the one that maps the average annual daily traffic (AADT) flows $\left\{F_{ij}\right\}$ to the counterfactual return on investments $\left\{R_{k\ell}^{\mathrm{cf}}\right\}$, which is
$R_{k\ell}^{\mathrm{cf}}=g_{k\ell}\left(\left\{F_{ij}\right\},\theta\right)$
for known functions $g_{k\ell}:\mathbb{R}_{+}^{n\left(n-1\right)}\times\Theta\rightarrow\mathbb{R}$ for $k,\ell=1,...,n$ locations in the US Interstate Highway network.
6.2.2 Measurement Error Model and Prior
For this application we can again apply the default approach from Section 4. For the empirical Bayes step we can use the baseline case from Section 4.2.1. There are no zeros so we only have to provide an estimate of the measurement error variance $\varsigma^{2}$. musunuru2019applications estimates that the measurement error variance of the logarithm of the average annual daily traffic (AADT) flows, which is exactly the data that allen2022welfare uses, is between 0.05 and 0.20. To obtain a lower bound on uncertainty, I will use a uniform measurement error variance of $0.05$.
With $\tilde{\varsigma}^{2}=0.05$, I use Equation (9) to find a prior variance of $\tilde{s}^{2}=0.101$. This results in the following estimated posterior distribution for the true traffic flow between country $i$ and $j$, $F_{ij}$, given its noisy version $\tilde{F}_{ij}$, for i,j=1,,ni,j=1,...,n:
$F_{ij}|\tilde{F}_{ij}\sim\mathrm{exp}\left\{\mathcal{N}\left(0.669\cdot\log\tilde{F}_{ij}+0.331\cdot\tilde{\mu}_{ij},0.033\right)\right\},$
where $\tilde{\mu}_{ij}$ is defined in Equation (8).
6.2.3 Results
The counterfactual question of interest is which links have the highest return on investment, and the authors of allen2022welfare report the top ten links. For exposition, I will focus my analysis on the three best performing links. Table 2 shows the $95\%$ intervals for the top three links based on Algorithm 2.
Point estimate Interval accounting for measurement error
Link 1 10.43 [8.69, 14.15]
Link 2 9.54 [7.31, 10.83]
Link 3 7.31 [6.78, 8.18]
Table 2: Uncertainty quantification for the return on investment for the three links from allen2022welfare with the highest return on investment. Returns on investment are reported as annualized decimal returns (10.43 means an annual 1043% return). Link 1 is Kingsport-Bristol (TN-VA) to Johnson City (TN), link 2 is Greensboro-High Point (NC) to Winston-Salem (NC) and link 3 is Rochester (NY) to Batavia (NY). The intervals based on measurement error report the 2.5th and 97.5th quantile of the estimated posterior distributions.
From a policy perspective it is of interest whether the ranking between these links can change due to measurement error. Therefore, Table 3 shows the $95\%$ intervals for the difference between link 1 and link 2, and the difference between link 2 and link 3.111111This simple exercise is intended purely for exposition. For a more formal treatment of inference on ranks, see mogstad2024inference. It follows that the rankings are generally robust against measurement error. Additional discussion and analyses can be found in Appendices E.2 and E.3.
Point estimate Interval accounting for measurement error
Link 1-Link 2 0.89 [0.38, 5.39]
Link 2-Link 3 2.23 [-0.05, 3.27]
Table 3: Uncertainty quantification for the differences in return on investment between the three links from allen2022welfare with the highest return on investment. Returns on investment are reported as annualized decimal returns (0.89 means an annual 89% return). Link 1 is Kingsport-Bristol (TN-VA) to Johnson City (TN), link 2 is Greensboro-High Point (NC) to Winston-Salem (NC) and link 3 is Rochester (NY) to Batavia (NY). The intervals based on measurement error report the 2.5th and 97.5th quantile of the estimated posterior distributions.
7 Conclusion
This paper develops an econometric framework for quantifying the impact of measurement error in a broad class of quantitative trade and spatial models. Unlike standard econometric models of measurement error, the counterfactual estimand in these models depends directly on the realized data rather than on the underlying distribution. I adopt an empirical Bayes approach to characterize uncertainty in counterfactual predictions and propose a default specification that can be easily implemented across a range of applications. Applying the framework to the settings in adao2017nonparametric and allen2022welfare, I find substantial uncertainty surrounding key economic outcomes. These results underscore the need to account for measurement error when using quantitative models to guide policy decisions.
Appendix
Appendix A Characterizing $g$ in Two Leading Classes of Models
Assumption 1 requires that, for a fixed counterfactual question and a fixed parameter value $\theta$, the counterfactual object of interest can be written as $\gamma=g\left(D,\theta\right)$, where $D$ collects the baseline inputs taken from the observed economy, and $\theta$ collects model parameters that are held fixed across baseline and counterfactual economies. This appendix discusses how to construct $g$ in two broad classes of quantitative models, namely invertible models and exact hat algebra models.
A.1 Invertible Models
Following redding2017quantitative, a model is invertible if, given $\theta$, the equilibrium conditions imply a one-to-one mapping between the baseline observables $D$ and a set of latent equilibrium objects $X\in\mathcal{X}\subseteq\mathbb{R}^{d_{X}}$, often called fundamentals. Formally, there exist maps $X=\mathcal{J}\left(D,\theta\right)$ and $D=\mathcal{E}\left(X,\theta\right)$, that are inverses of each other (up to standard normalizations such as a price level or numeraire). In quantitative trade and spatial models, $X$ often includes objects such as trade costs and productivity levels. For example, in the Armington model in Section 5, the fundamentals are $X=\left(\left\{\tau_{ij}\right\},\left\{\chi_{ij}\right\},\left\{\kappa_{i}\right\}\right)$.
Now consider a counterfactual experiment that specifies proportional changes to the fundamentals $X$. Denote these proportional changes by $X^{\mathrm{cf,prop}}\in\mathbb{R}^{d_{X}}$. For example in the Armington model in Section 5, the proportional changes are $X^{\mathrm{cf,prop}}=\left(\left\{1+0.1\cdot\mathbb{I}\left\{i\neq j\right\}\right\},\left\{1\right\},\left\{1\right\}\right)$. The goal is to compute the corresponding proportional changes to the observed data, $D^{\mathrm{cf,prop}}\in\mathbb{R}^{d_{D}}$, and then the scalar counterfactual object $\gamma$. The computational steps are:
  1. 1.
    “Invert” the model to recover the levels of the fundamentals $X$:
    $X=\mathcal{J}\left(D,\theta\right).$
  2. 2.
    Solve for counterfactual levels of the data $D\varodot D^{\mathrm{cf,prop}}$ and convert to proportional changes:
    $D\varodot D^{\mathrm{cf,prop}}=\mathcal{E}\left(X\varodot X^{\mathrm{cf,prop}},\theta\right)\quad\Rightarrow\quad D^{\mathrm{cf,prop}}=\mathcal{E}\left(X\varodot X^{\mathrm{cf,prop}},\theta\right)\oslash D,$
[…]
  3. 3.
    Compute the counterfactual object of interest:
    $\gamma=\Gamma\left(D,D^{\mathrm{cf,prop}},\theta\right).$
Composing these steps yields a function of $\left(D,\theta\right)$ alone:
$\gamma=\Gamma\left(D,\mathcal{E}\left(\mathcal{J}\left(D,\theta\right)\varodot X^{\mathrm{cf,prop}},\theta\right)\oslash D,\theta\right)\equiv g\left(D,\theta\right).$
A.2 Exact Hat Algebra Models
Exact hat algebra models (costinot2014trade) are models in which counterfactuals can be computed directly in proportional changes (“hats”) from baseline observables. In such models, there exist a mapping, given $\theta$, from the proportional changes in fundamentals and the observed data to the corresponding changes in the observed data,
$D^{\mathrm{cf,prop}}=\mathcal{H}\left(X^{\mathrm{cf,prop}},D,\theta\right).$
This mapping holds holds directly, without the intermediary step of backing out the levels of the fundamentals $X$. It follows that in this case
$\gamma=\Gamma\left(D,\mathcal{H}\left(X^{\mathrm{cf,prop}},D,\theta\right),\theta\right)\equiv g\left(D,\theta\right).$
The Armington model presented in Section 5 is an example of an exact hat algebra model. Since both invertible models and exact hat algebra models satisfy Assumption 1, they are essentially equivalent for the purposes of uncertainty quantification.
Appendix B Details for Relation to DingelTintelnot:2025
B.1 Details for Covariates-Based Approach
The specific form for $h_{ij}\left(\vartheta\right)$ in Equation (5) is
$h_{ij}\left(\vartheta\right)=\frac{w_{j}^{\varepsilon}\left(r_{i}^{\eta}\bar{\delta}_{ij}\right)^{-\varepsilon}}{\sum_{s,t}w_{t}^{\varepsilon}\left(r_{s}^{\eta}\bar{\delta}_{st}\right)^{-\varepsilon}},$
where $w_{j}$ denotes the wage in location $j$, $r_{i}$ denotes the land rent in location $i$, $\bar{\delta}_{ij}$ denotes the time component of commuting cost, $\varepsilon$ denotes the commuting elasticity, and $\eta$ denotes a Cobb-Douglas preference parameter. Since $\left\{\bar{\delta}_{ij}\right\}$ are known, it follows that $\vartheta=\left(\left\{\alpha_{i}^{\mathrm{orig}}\right\},\left\{\alpha_{j}^{\mathrm{dest}}\right\},\varepsilon\right)=\left(\left\{r_{i}^{-\eta\varepsilon}\right\},\left\{w_{j}^{\varepsilon}\right\},\varepsilon\right)$.
B.2 Using Matrix Approximation Techniques
In the truncated singular value decomposition approach in DingelTintelnot:2025, the recommendation is to use an approximated matrix instead of the matrix with noisy flows $\left\{\tilde{\ell}_{ij}\right\}$. To see this approach can be nested in my Bayesian framework by, defining $\tilde{\mathcal{L}}\equiv\left\{\tilde{\ell}_{ij}\right\}$ and $\mathcal{L}\equiv\left\{\ell_{ij}\right\}$, consider the following prior and measurement error:
$\begin{cases}\begin{array}[]{c}\mathrm{prior:}\\ \mathrm{measurement\>error:}\end{array}&\begin{array}[]{c}\pi^{\mathrm{prior}}\left(\mathcal{L}|A\right)=\delta_{A},\quad A\in\mathcal{A}=\left\{B:\mathrm{rank}\left(B\right)\leq\tau\right\}\\ \tilde{\mathcal{L}}=\mathcal{L}+\xi,\quad\xi_{ij}\overset{\mathrm{iid}}{\sim}\mathcal{N}\left(0,1\right)\end{array}.\end{cases}$ (14)
In this case, the empirical Bayes step solves:
$\tilde{A}$ $=\underset{A\in\mathcal{A}}{\arg\max}\int\exp\left(-\frac{1}{2}\sum_{i,j}\left(\tilde{\mathcal{L}}_{ij}-\mathcal{L}_{ij}\right)^{2}\right)\delta_{A}\left(\mathcal{L}\right)d\mathcal{L}$
$=\underset{A\in\mathcal{A}}{\arg\max}\exp\left(-\left\|\tilde{\mathcal{L}}-A\right\|_{F}^{2}\right),$
where $\left\|\cdot\right\|_{F}$ denotes the Frobenius norm. This maximization problem is equivalent to projecting the noisy flows onto the space of matrices that have a rank no larger than $\tau$, which, by the Eckart–Young–Mirsky theorem, is solved by the truncated singular value decomposition. This yields the estimated prior and measurement error model
$\begin{cases}\begin{array}[]{c}\mathrm{prior:}\\ \mathrm{measurement\>error:}\end{array}&\begin{array}[]{c}\pi^{\mathrm{prior}}\left(\mathcal{L}|\tilde{A}\right)=\delta_{\tilde{A}},\quad A\in\mathcal{A}=\left\{B:\mathrm{rank}\left(B\right)\leq\tau\right\}\\ \tilde{\mathcal{L}}=\mathcal{L}+\xi,\quad\xi_{ij}\overset{\mathrm{iid}}{\sim}\mathcal{N}\left(0,1\right)\end{array}.\end{cases}$
Using Bayes’ rule we can then find the estimated posterior for the true flows
$\pi^{\mathrm{post}}\left(\mathcal{L}|\tilde{\mathcal{L}},\tilde{A}\right)=\delta_{\tilde{A}}$.
So the posterior is a point mass at the counterfactual prediction that uses the approximated matrix $\tilde{A}$. It follows that the truncated singular value decomposition approach is a special case of Algorithm 1 by choosing the prior and measurement error model as in Equation (14).
Appendix C Details for Estimation Error
The counterfactual prediction of interest will typically depend on a structural parameter $\theta$. It is common in applied work to plug in a fixed value for the structural parameter taken from the literature or obtained through data-driven methods, thus ignoring the uncertainty associated with the estimation process. I will discuss two different approaches to dealing with estimation error.
C.1 Frequentist Approach to Dealing with Estimation Error
Let $\tilde{\theta}$ denote the estimator of the estimand $\theta$. This estimand is usually a function of the distribution of the data $\mathcal{P}_{D}$. This implies that, to address measurement error affecting the structural parameter, one can apply the frequentist measurement error techniques discussed in Section 3.4.1 to find a bias-corrected estimate, though the resulting correction will not admit a Bayesian interpretation. For example, one may use repeated measurements or instrumental variables to construct consistent estimators.
C.2 Bayesian Approach to Dealing with Estimation Error
Alternatively, one can take a Bayesian or quasi-Bayesian approach and assume that the posterior or quasi-posterior distribution of the true structural parameter $\theta$ given the data $D$ is approximately normal. [^12] Formally, this normality could follow from assumptions on the underlying data generating process such that a Bernstein-von Mises type result holds (van2000asymptotic). In that case the influence of the prior distribution $\pi\left(\theta\right)$ becomes negligible and the posterior distribution approximately equals a normal distribution centered at the maximum likelihood estimator. In sanders2025new I engage further with structural estimation in quantitative trade and spatial models. Specifically, we have the estimation error posterior (with superscript “post,ee”)
$\pi^{\mathrm{post,ee}}\left(\theta|D\right)\approx\mathcal{N}\left(\tilde{\theta}\left(D\right),\tilde{\Sigma}\left(D\right)\right),$ (15)
[…]
We can then generate draws from the posterior distribution of $\theta$ given DD. [^14] Note that this assumption is on the structural parameter $\theta$, and not on the fundamentals as discussed in Appendix A. One could additionally use a degenerate posterior $\pi^{\mathrm{post,ee}}\left(X|D\right)=\delta_{X}$ on these fundamentals, since they are parameters that are linked deterministically to the data $D$, and hence there is no estimation error. For each of these draws, we can calculate the corresponding value of the counterfactual object of interest using the relationship $\gamma=g\left(D,\theta\right)$. This allows us to find the posterior distribution of $\gamma$ given the true data, $\pi^{\mathrm{post,ee}}\left(\gamma|D\right)$.
C.2.1 Combining Measurement Error and Estimation Error
The object of interest is a function of the true data and the structural parameter. It follows that we must consider estimation error, the direct effect of mismeasurement, and the indirect effect of mismeasurement through the estimation procedure. Our goal is to quantify uncertainty about $\gamma$ when we observe $\tilde{D}$ by accounting for these various sources of uncertainty.
[…]
For uncertainty quantification, I recommend to report an interval $\mathcal{C}$ to which, in posterior expectation over $D$, the posterior $\pi^{\mathrm{post,ee}}\left(\gamma|D\right)$ assigns probability $1-\alpha$:
$\mathbb{E}_{\pi^{\mathrm{post,me}}}\left[Pr_{\pi^{\mathrm{post,ee}}}\left\{\gamma\in\mathcal{C}|D\right\}|\tilde{D},\tilde{\vartheta}\right]\geq 1-\alpha.$
[…]
Algorithm 4 Uncertainty quantification about $\gamma=g\left(D,\theta\right)$
  1. 1.
    Input: prior $\pi^{\mathrm{prior}}\left(D|\vartheta\right)$, measurement error model $\pi^{\mathrm{me}}\left(\tilde{D}|D,\vartheta\right)$, quasi-posterior $\pi^{\mathrm{post,ee}}\left(\theta|D\right)$, noisy data $\tilde{D}$, number of bootstrap draws $B$, coverage level $1-\alpha$ (choose $B$ and $\alpha$ such that $\alpha/2\cdot B\in\mathbb{N}$).
  2. 2.
    Empirical Bayes estimation step: $\tilde{\vartheta}=\underset{\vartheta}{\arg\max}\int\pi^{\mathrm{me}}\left(\tilde{D}|D,\vartheta\right)\pi^{\mathrm{prior}}\left(D|\vartheta\right)dD.$
  3. 3.
[…]
  4. 4.
    For b=1,,Bb=1,...,B,
    1. (a)
[…]
    2. (b)
      Draw $\theta_{b}\sim\pi^{\mathrm{post,ee}}\left(\theta|D_{b}\right)$.
    3. (c)
      Compute $\gamma_{b}=g\left(D_{b},\theta_{b}\right).$
  5. 5.
    Sort $\left\{\gamma_{b}\right\}_{b=1}^{B}$ to obtain $\left\{\gamma^{(b)}\right\}_{b=1}^{B}$ with $\gamma^{(1)}\leq\gamma^{(2)}\leq...\leq\gamma^{(B)}$.
  6. 6.
    Report $\left[\gamma^{(\alpha/2\cdot B)},\gamma^{(\left(1-\alpha/2\right)\cdot B)}\right].$
As in Section 3.3, a natural point estimator for the structural parameter is the median of the estimated posterior of the structural parameter given the noisy data,
$\pi^{\mathrm{post}}\left(\theta|\tilde{D},\tilde{\vartheta}\right)=\int\pi^{\mathrm{post,ee}}\left(\theta|D\right)\pi^{\mathrm{post,me}}\left(D|\tilde{D},\tilde{\vartheta}\right)dD,$
and a natural point estimator for the counterfactual prediction is the median of the estimated posterior,
[…]
These posterior medians can be calculated using the draws obtained in steps 4b and 4c in Algorithm 4, respectively.
I illustrate this procedure in Appendix D.3 by applying it to the application in adao2017nonparametric.
C.2.2 Frequentist Consistency
A caveat of the proposed Bayesian approach to dealing with estimation error is that it will not guarantee frequentist consistency of the estimator. To see this, consider the stylized model in which the independent variable in a simple regression is measured with error:
$\begin{cases}\begin{array}[]{c}\mathrm{regression:}\\ \mathrm{measurement\ error:}\\ \mathrm{prior:}\end{array}&\begin{array}[]{c}\theta=\frac{\mathrm{Cov}\left(Y_{i},X_{i}\right)}{\mathrm{Var}\left(X_{i}\right)}\\ \tilde{X}_{i}=X_{i}+\varepsilon_{i}\\ X_{i}=\beta Z_{i}+\nu_{i}\end{array},\end{cases}$
where $\varepsilon_{i}$, $Z_{i}$ and $\nu_{i}$ are mean-zero normal random variables. In this setting, the IV estimator $\frac{\mathrm{\widehat{Cov}}\left(Y_{i},Z_{i}\right)}{\mathrm{\widehat{Cov}}\left(\tilde{X}_{i},Z_{i}\right)}$ is consistent provided an exogeneity and relevance condition hold.
Alternatively, using the proposed empirical Bayesian approach, we would estimate the hyperparameter $\beta$ by noting that
$\tilde{X}_{i}=\beta Z_{i}+\nu_{i}+\varepsilon_{i},$
and we would use the empirical Bayes estimator
$\tilde{\beta}=\frac{\mathrm{\widehat{Cov}}\left(\tilde{X_{i}},Z_{i}\right)}{\mathrm{\widehat{Var}}\left(Z_{i}\right)}\overset{p}{\rightarrow}\beta+\frac{\mathrm{\mathrm{Cov}}\left(\nu_{i}+\varepsilon_{i},Z_{i}\right)}{\mathrm{\mathrm{Var}}\left(Z_{i}\right)}.$
[…]
$X_{i}^{*}$ $=w_{\mathrm{po}}\cdot\tilde{X}_{i}+\left(1-w_{\mathrm{po}}\right)\cdot\tilde{\beta}Z_{i}$
$\overset{p}{\rightarrow}w_{\mathrm{po}}\cdot\tilde{X}_{i}+\left(1-w_{\mathrm{po}}\right)\cdot\beta Z_{i}+\left(1-w_{\mathrm{po}}\right)\frac{\mathrm{\mathrm{Cov}}\left(\nu_{i}+\varepsilon_{i},Z_{i}\right)}{\mathrm{\mathrm{Var}}\left(Z_{i}\right)}Z_{i}$
$=X_{i}+\underbrace{w_{\mathrm{po}}\cdot\varepsilon_{i}-\left(1-w_{\mathrm{po}}\right)\cdot\nu_{i}+\left(1-w_{\mathrm{po}}\right)\frac{\mathrm{\mathrm{Cov}}\left(\nu_{i}+\varepsilon_{i},Z_{i}\right)}{\mathrm{\mathrm{Var}}\left(Z_{i}\right)}Z_{i}}_{\equiv p_{i}}.$
Here, $p_{i}$ captures the deviation of the posterior draw from the truth, which has mean zero but non-zero variance. The probability limit of the regression coefficient based on posterior draws then is
$\frac{\mathrm{Cov}\left(Y_{i},X_{i}+p_{i}\right)}{\mathrm{Var}\left(X_{i}+p_{i}\right)},$
which does not generally equal $\theta=\frac{\mathrm{Cov}\left(Y_{i},X_{i}\right)}{\mathrm{Var}\left(X_{i}\right)}$.
Appendix D Details for Application adao2017nonparametric
D.1 Model Details
In the empirical application of adao2017nonparametric, the authors investigate the effects of China joining the WTO, the so-called China shock. Going forward, $Q_{i,t}$ denotes the factor endowment of country $i$ in period $t$, $\tau_{ij,t}$ denotes the trade cost between country $i$ and $j$ in period $t$, $\lambda_{ij,t}$ denotes the expenditure share from country $i$ in country $j$ in period $t$, $Y_{i,t}$ denotes the income of country $i$ in period $t$, and $P_{i,t}$ denotes the factor price of country $i$ in period $t$. Furthermore, $\rho_{i,t}$ denotes the difference between aggregated gross expenditure and gross production in country $i$ in period $t$, which is assumed to stay constant for different counterfactuals. Lastly, $\varepsilon$ denotes the trade elasticity and $\chi_{i}\left(\cdot\right)$ denotes the factor demand system of country $i$.
In adao2017nonparametric, two demand systems are considered, normal CES and “Mixed CES”. I will focus on normal CES, so that
$\lambda_{ij,t}=\chi_{i}\left(\left\{\delta_{ij,t}\right\}\right)=\frac{\exp\left\{\delta_{ij,t}\right\}}{1+\sum_{\ell>1}\exp\left\{\delta_{i\ell,t}\right\}},$
for $\delta_{ij,t}$ some transformation of factor prices. The function $\chi_{i}^{-1}\left(\cdot\right)$ then maps the observed expenditures shares to values of this transformation. The structural parameter $\varepsilon$ is estimated by assuming a model on the unobserved trade costs $\left\{\tau_{ij,t}\right\}$, and is estimated using GMM with as an input the expenditure shares $\left\{\lambda_{ij,t}\right\}$.
The counterfactual question of interest is what the change in China’s welfare is due to joining the WTO. This question is modeled by choosing the counterfactual proportional changes in trade costs, $\left\{\tau_{ij,t}^{\mathrm{cf,prop}}\right\}$, such that Chinese trade costs are brought back to their 1995 levels:
$\tau_{ij,t}^{\mathrm{cf,prop}}$ $=\frac{\tau_{ij,95}}{\tau_{ij,t}},\quad\mathrm{if}\>i\>\mathrm{or}\>j\>\mathrm{is}\>\mathrm{China,}$
$\tau_{ij,t}^{\mathrm{cf,prop}}$ $=1,\quad\mathrm{otherwise.}$
Welfare is then defined as the percentage change in income that the representative agent in China would be indifferent about accepting instead of the counterfactual change in trade costs from $\left\{\tau_{ij,t}\right\}$ to $\left\{\tau_{ij,t}^{\mathrm{cf,prop}}\tau_{ij,t}\right\}$. These proportional changes in China’s welfare $\left\{W_{\mathrm{China},t}^{\mathrm{cf,prop}}\right\}$ can be obtained from first solving for $\left\{P_{i,t}^{\mathrm{cf,prop}}\right\}$ using the system of equations
$\sum_{j}\frac{\exp\left\{\chi_{i}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)-\varepsilon\log\left(P_{i,t}^{\mathrm{cf,prop}}\tau_{ij,t}^{\mathrm{cf,prop}}\right)\right\}}{1+\sum_{\ell>1}\exp\left\{\chi_{\ell}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)-\varepsilon\log\left(P_{\ell,t}^{\mathrm{cf,prop}}\tau_{\ell j,t}^{\mathrm{cf,prop}}\right)\right\}}\left\{P_{j,t}^{\mathrm{cf,prop}}Y_{j,t}+\rho_{j,t}\right\}=P_{i,t}^{\mathrm{cf,prop}}Y_{i,t},$
and then using
$W_{i,t}^{\mathrm{cf,prop}}=100\cdot\left(P_{i,t}^{\mathrm{cf,prop}}\frac{\sum_{\ell}\left[\chi_{\ell}^{-1}\left(\left\{\lambda_{ij,t}\right\}\right)\right]^{-\varepsilon}}{\sum_{\ell}\left[P_{\ell,t}^{\mathrm{cf,prop}}\tau_{i\ell,t}^{\mathrm{cf,prop}}\left(\left\{\lambda_{ij,t}\right\}\right)\right]^{-\varepsilon}}-1\right).$
D.2 Calibration Procedure and Computational Details
The default approach from Section 4 can be applied. For the empirical Bayes step, the calibration of $\vartheta$, we can use the mirror trade data setting from Section 4.2.2.
In preprocessing the mirror trade dataset from linsi2023problem I made some additional assumptions. Firstly, I only consider data from the period that is considered in adao2017nonparametric. Secondly, I only consider trade flows between countries that the authors of that paper consider. This amounts to aggregating Belgium and Luxembourg, and Estonia and Latvia. All the remaining countries I aggregate to “Rest of World”. Thirdly, when only one of the mirror trade flows is reported, I interpret this as zero measurement error by setting the unknown mirror trade flow equal to the observed one. Relatedly, when both mirror trade flows are not reported, I interpret this as there being no trade, and when one trade flow is zero and the other is substantially larger than zero, I set the zero trade flow equal to the non-zero one. Lastly, I follow adao2017nonparametric by setting zero trade flows to 0.0025 (million USD). There are however only a handful of zeros due to the aggregation into “Rest of World”.
When estimating the prior distribution of the true underlying trade flows, I use the distance dataset from mayer2011notes. For the distance between countries and the “Rest of World”, I take the average of the distances to all other countries that are considered in adao2017nonparametric.
An important consideration is that there is a substantial difference between the trade flows used in adao2017nonparametric, which come from the World Input Output Dataset (WIOD), and the mirror trade flows from linsi2023problem, which are based on the IMF Direction of Trade Statistics dataset. To overcome this discrepancy, I scale the mirror trade data to make them comparable to the trade flows from WIOD. I set $\tilde{F}_{ij,t}^{1,\mathrm{ACD}}=\tilde{F}_{ij,t}^{\mathrm{ACD}}$ and $\tilde{F}_{ij,t}^{2,\mathrm{ACD}}=\tilde{F}_{ij,t}^{2}\cdot\tilde{F}_{ij,t}^{\mathrm{ACD}}/\tilde{F}_{ij,t}^{1}$, for $\tilde{F}_{ij,t}^{\mathrm{ACD}}$ the noisy trade flow as used in adao2017nonparametric. There were also some trade flows in the mirror trade dataset that reported zeros but had a large trade flow in the WIOD. For these trade flows, I set the zero mirror trade data entries equal to the positive WIOD entry.
D.3 Supplementary Analyses
D.3.1 Winsorized Measurement Error Variances
The distribution of measurement error variances has a heavy right tail, with the noisiest bilateral trade flow the one from Mexico to Australia with a measurement error variance of 1.42. One might be worried that this heavy tail drives the sensitivity to mismeasurement. Figure 3 replicates Figure 2 but now winsorizing the measurement error variances at 0.2, but keeping the posterior variances constant. This amounts to winsorizing 27% of the trade flows. There are no substantial differences between Figures 3 and 2.
D.3.2 Accounting for Estimation Error and Measurement Error Simultaneously
As mentioned in Section 6.1, the authors of adao2017nonparametric estimate the trade elasticity $\varepsilon$ using GMM. Figure 4 replicates Figure 2 but adds intervals obtained using Algorithm 4. Interestingly, the upper bounds increase substantially, while the lower bounds remain close to the lower bounds of the intervals that account only for measurement error.
D.3.3 Testing Normality Assumption and Gravity Model for the Prior
As outlined in Remark 1, we can check how reasonable the normality assumption is by comparing the histogram of the normalized residuals with the probability density function of a standardized normal distribution. The result can be found in Figure 5. It follows that the normality assumption seems reasonable.
Concerning the gravity model, restricting attention to the year 2011, the regression for the prior mean in Equation (17) has an adjusted R-squared of 0.95, and the coefficient on log distance is -0.277 with a t-statistic of 3.346. Furthermore, Figure 6 follows allen201813 by plotting a linear and nonparametric fit of log trade flows against log distance, after partitioning out the origin and destination fixed effects. Together, the high adjusted R-squared and the good performance of the linear fit imply that the gravity model is a reasonable choice for this setting.
Appendix E Details for Application allen2022welfare
E.1 Model Details
In the empirical application of allen2022welfare, the authors investigate what the returns on investment are of all the highway segments of the US Interstate Highway network. Going forward, $\bar{L}$ denotes aggregate labor endowment, $\bar{Y}$ denotes total income in the economy, $Q_{i}$ denotes the productivity of location $i$, $A_{i}$ captures the level of amenities in location $i$, $\tau_{ij}$ denotes the travel cost between locations $i$ and $j$, $F_{ij}$ denotes the traffic flow between locations $i$ and $j$, $y_{i}$ denotes total income of location $i$ as a share of the total income in the economy, $\ell_{i}$ denotes the total labor in location $i$ as a share of the aggregate labor endowment, and $\chi$ captures the (inverse of) the welfare of the economy. The parameter vector is $\theta=\left(\alpha,\beta,\gamma,\nu\right)$, where $\alpha$ and $\beta$ control the strength of the productivity and amenity externalities respectively, $\gamma$ is the shape parameter of the Fréchet distributed idiosyncratic productivity shocks, and $\nu$ governs the strength of traffic congestion.
It is shown in the paper that we can uniquely recover $\left(\left\{y_{i}^{\mathrm{cf,prop}}\right\},\left\{\ell_{i}^{\mathrm{cf,prop}}\right\},\chi^{\mathrm{cf,prop}}\right)$ given any change in the underlying infrastructure network $\left\{\tau_{ij}^{\mathrm{cf,prop}}\right\}$ and baseline economic activity $\left\{y_{i}\bar{Y}\right\}$, using the system of equations
$\left(y_{i}^{\mathrm{cf,prop}}\right)^{\frac{1+\nu+\gamma}{1+\nu}}\left(\ell_{i}^{\mathrm{cf,prop}}\right)^{\frac{-\theta\left(1+\alpha+\nu\left(\alpha+\beta\right)\right)}{1+\nu}}$
$=\chi^{\mathrm{cf,prop}}\left(\frac{y_{i}\bar{Y}}{y_{i}\bar{Y}+\sum_{k}F_{ik}}\right)\left(y_{i}^{\mathrm{cf,prop}}\right)^{\frac{1+\nu+\gamma}{1+\nu}}\left(\ell_{i}^{\mathrm{cf,prop}}\right)^{\frac{\gamma\left(\beta-1\right)}{1+\nu}}$
$\qquad+\sum_{j}\left(\frac{F_{ij}}{y_{i}\bar{Y}+\sum_{k}F_{ik}}\right)\left(\tau_{ij}^{\mathrm{cf,prop}}\right)^{\frac{-\gamma}{1+\nu}}\left(y_{j}^{\mathrm{cf,prop}}\right)^{\frac{1+\gamma}{1+\nu}}\left(\ell_{j}^{\mathrm{cf,prop}}\right)^{\frac{-\gamma\left(1+\alpha\right)}{1+\nu}}$
$\left(y_{i}^{\mathrm{cf,prop}}\right)^{\frac{-\gamma+\nu}{1+\nu}}\left(\ell_{i}^{\mathrm{cf,prop}}\right)^{\frac{\gamma\left(1-\beta-\nu\left(\alpha+\beta\right)\right)}{1+\nu}}$
$=\chi^{\mathrm{cf,prop}}\left(\frac{y_{i}\bar{Y}}{y_{i}\bar{Y}+\sum_{k}F_{ki}}\right)\left(y_{i}^{\mathrm{cf,prop}}\right)^{\frac{-\gamma+\nu}{1+\nu}}\left(\ell_{i}^{\mathrm{cf,prop}}\right)^{\frac{\gamma\left(\alpha+1\right)}{1+\nu}}$
$\qquad+\sum_{j}\left(\frac{F_{ji}}{y_{i}\bar{Y}+\sum_{k}F_{ki}}\right)\left(\tau_{ij}^{\mathrm{cf,prop}}\right)^{\frac{-\gamma}{1+\nu}}\left(y_{j}^{\mathrm{cf,prop}}\right)^{\frac{-\gamma}{1+\nu}}\left(\ell_{j}^{\mathrm{cf,prop}}\right)^{\frac{\gamma\left(1-\beta\right)}{1+\nu}}.$
Having obtained $\chi^{\mathrm{cf,prop}}$, the proportional counterfactual change in welfare is then calculated using
$W^{\mathrm{cf,prop}}=\frac{\left(\chi^{\mathrm{cf,prop}}\right)^{1/\gamma}}{\bar{L}^{\alpha+\beta}}.$
E.2 Calibration Procedure and Computational Details
The default approach from Section 4 can be applied. For the empirical Bayes step, the calibration of $\vartheta$, we can use the baseline case with domain knowledge from Section 4.2.1.
When I run the code from allen2022welfare, the returns of investment for the links systematically differ slightly from the ones in the paper. I scale my estimates so that the unperturbed estimates align with the ones in the paper.
E.3 Supplementary Analyses
E.3.1 Probability that Rankings are Reversed
We can learn more from the posterior distributions than just intervals. It might be of interest what the expected probability is that the ranking of the three links are reversed. The expected probability that the ranking between link 1 and link 2 is reversed is 0.000, and the expected probability that the ranking between link 2 and link 3 is 0.030.
E.3.2 Testing Normality Assumption and Gravity Model for the Prior
We can again check the reasonableness of the normality assumption as per Remark 1. The result can be found in Figure 7, and it follows that the normality assumption is less reasonable compared to the setting of adao2017nonparametric.
Concerning the gravity model, the regression for the prior mean in Equation (8) has an adjusted R-squared of $0.9995$, and the coefficient on log distance is $1.003$ with a t-statistic of $1138$. It follows that log distance is an important driver of log traffic flows, but not in a negative way as is common in gravity models. Furthermore, Figure 8 follows allen201813 by plotting a linear and nonparametric fit of log traffic flows against log distance, after partitioning out the origin and destination fixed effects. Together, the high adjusted R-squared and the good performance of the linear fit imply that the gravity model is a reasonable choice for this setting.
Appendix F Misspecification of the Measurement Error Model and Prior
We are interested in the potential effects of misspecification of the measurement error model or prior. Specifically, focusing on the widely applicable default approach from Section 4, we would like to know how the quantiles of the posterior distribution of the counterfactual object of interest given the noisy flows change when the assumptions of a normal measurement error model or a normal prior do not hold. Suppose for exposition that there are no zeros and the hyperparameters $\vartheta$ are known, so that we can obtain the posterior distribution $\pi^{\mathrm{post}}\left(\gamma|\left\{\log\tilde{F}_{ij}\right\}\right)$.
F.1 Measurement Error Model
Let $L\left(\left\{\log F_{ij}\right\}\right)=\pi^{\mathrm{me}}\left(\left\{\log\tilde{F}_{ij}\right\}|\left\{\log F_{ij}\right\}\right)$ denote the likelihood function of the noisy log flows $\left\{\log\tilde{F}_{ij}\right\}$ given the true log flows $\left\{\log F_{ij}\right\}$. For a given $c\geq 1$, define a density-ratio class of distributions to be the set of all conditional distributions for $\left\{\log\tilde{F}_{ij}\right\}$ with pdf $p$ such that
$p\in\mathcal{R}_{c}=\left\{p\in P:\frac{1}{c}\cdot L\left(x\right)\leq p\left(x\right)\leq c\cdot L\left(x\right)\ \forall x\in\mathbb{R}^{n\left(n+1\right)}\right\},$
for $P$ the set of all pdfs.
For uncertainty quantification, we are interested in the quantiles of the posterior distribution $\pi^{\mathrm{post}}\left(h\left(\left\{\log F_{ij}\right\}\right)|\left\{\log\tilde{F}_{ij}\right\}\right)$ for a generic function $h\left(\cdot\right)$. Denote the $\alpha$-th posterior quantile based on likelihood $p$ by $Q_{\pi,p,h}\left(\alpha\right)$.
Proposition 1.
We have:
$\underset{p\in\mathcal{R}_{c}}{\sup}Q_{\pi,p,h}\left(\alpha\right)$ $=Q_{\pi,L,h}\left(\frac{\alpha c^{2}}{1-\alpha+\alpha c^{2}}\right)$
$\underset{p\in\mathcal{R}_{c}}{\inf}Q_{\pi,p,h}\left(\alpha\right)$ $=Q_{\pi,L,h}\left(\frac{\alpha}{\alpha+\left(1-\alpha\right)c^{2}}\right).$
So instead of reporting the interval
$\left[Q_{\pi,L,h}\left(\alpha/2\right),Q_{\pi,L,h}\left(1-\alpha/2\right)\right]$
one could report the robust interval
$\left[Q_{\pi,L,h}\left(\frac{\alpha}{\alpha+\left(2-\alpha\right)c^{2}}\right),Q_{\pi,L,h}\left(\frac{\left(2-\alpha\right)c^{2}}{\alpha+\left(2-\alpha\right)c^{2}}\right)\right].$
For example for $\alpha=0.05$ and $c=1.5$, we would consider the $1.1\%$-quantile and the $98.9\%$-quantile, instead of the $2.5\%$-quantile and the $97.5\%$-quantile, respectively.
The result in Proposition 1 follows from noting that
$\alpha=\int_{-\infty}^{q}\pi^{\mathrm{post}}\left(h\left(x\right)|\tilde{x}\right)dh\left(x\right)=\int_{x\in h^{-1}\left(\left[-\infty,q\right]\right)}\pi^{\mathrm{post}}\left(x|\tilde{x}\right)dx$
$\Rightarrow$ $\int_{x\in h^{-1}\left(\left[-\infty,q\right]\right)}p\left(x\right)\pi^{\mathrm{prior}}\left(x\right)dx=\frac{\alpha}{1-\alpha}\int_{x\not\in h^{-1}\left(\left[-\infty,q\right]\right)}p\left(x\right)\pi^{\mathrm{prior}}\left(x\right)dx.$
Focusing on the upper bound, it follows that we want to choose $p\left(x\right)$ on the left-hand side as small as possible and $p\left(x\right)$ on the right-hand side as large as possible for all $x$:
$\frac{1}{c}\int_{x\in h^{-1}\left(\left[-\infty,q_{\mathrm{sup}}^{*}\right]\right)}L\left(x\right)\pi^{\mathrm{prior}}\left(x\right)dx=\frac{\alpha}{1-\alpha}c\int_{x\not\in h^{-1}\left(\left[-\infty,q_{\mathrm{sup}}^{*}\right]\right)}L\left(x\right)\pi^{\mathrm{prior}}\left(x\right)dx$
$\Rightarrow$ $\int_{-\infty}^{q_{\mathrm{sup}}^{*}}\pi^{\mathrm{post}}\left(h\left(x\right)|\left\{\log\tilde{F}_{ij}\right\}\right)dh\left(x\right)=\frac{\alpha c^{2}}{1-\alpha+\alpha c^{2}}.$
F.2 Prior
Note that the likelihood $L$ and the prior $\pi^{\mathrm{prior}}$ enter the posterior in exactly the same way, so we can interpret the procedure in the previous subsection also as sensitivity analysis with respect to the prior.
Appendix G Calibration with Mirror Trade Data
G.1 Model
I use the mirror trade dataset from linsi2023problem. This dataset has two estimates of each bilateral trade flow, both as reported by the exporter and as by the importer. I interpret this as observing two independent noisy observations per time period for each bilateral trade flow: $\left\{\left\{\tilde{F}_{ij,t}^{1},\tilde{F}_{ij,t}^{2}\right\}_{t=1}^{T}\right\}_{i\neq j}$. It is helpful to rewrite the model:
$\begin{cases}\begin{array}[]{c}\mathrm{true\ zeros:}\\ \mathrm{spurious\ zeros:}\\ \mathrm{prior:}\\ \\ \\ \mathrm{likelihood:}\\ \\ \\ \\ \\ \end{array}&\begin{array}[]{c}P_{ij,t}\sim\mathrm{Bern}\left(p_{ij}\right)\\ B_{ij,t}^{1},B_{ij,t}^{2}\sim\mathrm{Bern}\left(b_{ij}\right)\\ F_{ij,t}\sim P_{ij,t}\cdot\delta_{0}+\left(1-P_{ij,t}\right)\cdot e^{\mu_{ij,t}}\cdot e^{\eta_{ij,t}}\\ \mu_{ij,t}=\beta_{t}\log\mathrm{dist}_{ij}+\alpha_{i,t}^{\mathrm{orig}}+\alpha_{j,t}^{\mathrm{dest}}\\ \eta_{ij,t}\sim\mathcal{N}\left(0,s_{ij}^{2}\right)\\ \tilde{F}_{ij,t}^{1}|F_{ij,t}\sim\delta_{0}\cdot\mathbb{I}\left\{F_{ij,t}=0\right\}\\ +\left[B_{ij,t}^{1}\cdot\delta_{0}+\left(1-B_{ij,t}^{1}\right)\cdot F_{ij,t}\cdot e^{\varepsilon_{ij,t}^{1}}\right]\cdot\mathbb{I}\left\{F_{ij,t}>0\right\}\\ \tilde{F}_{ij,t}^{2}|F_{ij,t}\sim\delta_{0}\cdot\mathbb{I}\left\{F_{ij,t}=0\right\}\\ +\left[B_{ij,t}^{2}\cdot\delta_{0}+\left(1-B_{ij,t}^{2}\right)\cdot F_{ij,t}\cdot e^{\varepsilon_{ij,t}^{2}}\right]\cdot\mathbb{I}\left\{F_{ij,t}>0\right\}\\ \varepsilon_{ij,t}^{1},\varepsilon_{ij,t}^{2}\sim\mathcal{N}\left(0,\varsigma_{ij}^{2}\right)\end{array}.\end{cases}$
G.2 Bernoulli Parameters
For a given bilateral trade flow from $i$ to $j$ in period $t$, we can compute the ex-ante probability of observing a certain number of zeros:
$Pr\left\{\mathrm{two\ observed\ zeros}\right\}$ $=p_{ij}+\left(1-p_{ij}\right)\cdot b_{ij}^{2}$
$Pr\left\{\mathrm{one\ observed\ zero}\right\}$ $=2\cdot\left(1-p_{ij}\right)\cdot\left(1-b_{ij}\right)\cdot b_{ij}$
$Pr\left\{\mathrm{no\ observed\ zeros}\right\}$ $=\left(1-p_{ij}\right)\cdot\left(1-b_{ij}\right)^{2}.$
We can use the time variation to identify the probabilities on the left-hand side:
$\tilde{z}_{ij,2}$ $=\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}\left\{\tilde{F}_{ij,t}^{1}=0,\tilde{F}_{ij,t}^{2}=0\right\}$
$\tilde{z}_{ij,1}$ $=\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}\left\{\tilde{F}_{ij,t}^{1}=0,\tilde{F}_{ij,t}^{2}>0\mathrm{\ or\ }\tilde{F}_{ij,t}^{1}>0,\tilde{F}_{ij,t}^{2}=0\right\}$
$\tilde{z}_{ij,0}$ $=\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}\left\{\tilde{F}_{ij,t}^{1}>0,\tilde{F}_{ij,t}^{2}>0\right\}.$
A large $T$ is not required for identification, but it improves precision. When $\tilde{z}_{ij,2},\tilde{z}_{ij,1},\tilde{z}_{ij,0}\in\left(0,1\right)$, we can back out the estimated probability of a true zero $\tilde{p}_{ij}$ and the estimated probability of a spurious zero $\tilde{b}_{ij}$ by solving
$\tilde{z}_{ij,2}$ $=\tilde{p}_{ij}+\left(1-\tilde{p}_{ij}\right)\cdot\tilde{b}_{ij}^{2}$
$\tilde{z}_{ij,1}$ $=2\cdot\left(1-\tilde{p}_{ij}\right)\cdot\left(1-\tilde{b}_{ij}\right)\cdot\tilde{b}_{ij}$
$\tilde{z}_{ij,0}$ $=\left(1-\tilde{p}_{ij}\right)\cdot\left(1-\tilde{b}_{ij}\right)^{2}.$
The solutions are
$\tilde{p}_{ij}=\max\left\{1-\frac{\left(\tilde{z}_{ij,1}+2\tilde{z}_{ij,0}\right)^{2}}{4\tilde{z}_{ij,0}},0\right\},\quad\tilde{b}_{ij}=\frac{\tilde{z}_{ij,1}}{\tilde{z}_{ij,1}+2\tilde{z}_{ij,0}}.$
I separately consider the possible cases where the estimated probabilities $\left(\tilde{z}_{ij,2},\tilde{z}_{ij,1},\tilde{z}_{ij,0}\right)$ are not all strictly between 0 and 1:
  1. 1.
    $\tilde{z}_{ij,2}=1,\tilde{z}_{ij,1}=0,\tilde{z}_{ij,0}=0$: In this case we observe only zeros so I set the estimated probability of a true zero $\tilde{p}_{ij}$ to $1$, which makes the estimated probability of a spurious zero $\tilde{b}_{ij}$ irrelevant.
  2. 2.
    $\tilde{z}_{ij,2}=0,\tilde{z}_{ij,1}=1,\tilde{z}_{ij,0}=0$: In this case one country always reports a positive flow and the other reports a zero flow. In this case I set the estimated probability of a true zero $\tilde{p}_{ij}$ to 0, and the estimated probability of a spurious zero $\tilde{b}_{ij}$ to $0.5$.
  3. 3.
    $\tilde{z}_{ij,2}=0,\tilde{z}_{ij,1}=0,\tilde{z}_{ij,0}=1$: In this case all reported flows are positive, so I set both the estimated probability of a true zero $\tilde{p}_{ij}$ and the estimated probability of a spurious zero $\tilde{b}_{ij}$ to 0.
  4. 4.
    $\tilde{z}_{ij,2}\in\left(0,1\right),\tilde{z}_{ij,1}\in\left(0,1\right),\tilde{z}_{ij,0}=0$: In this case there are no years with two reported positive flows. In this case I set the estimated probability of a true zero $\tilde{p}_{ij}$ to $\tilde{z}_{ij,2}$, and the estimated probability of a spurious zero $\tilde{b}_{ij}$ to $\tilde{z}_{ij,1}$.
  5. 5.
    $\tilde{z}_{ij,2}\in\left(0,1\right),\tilde{z}_{ij,1}=0,\tilde{z}_{ij,0}\in\left(0,1\right)$: In this case some years have two zeros and other years have two positive flows. In this case I set the estimated probability of a true zero $\tilde{p}_{ij}$ to $\tilde{z}_{ij,2}$, and the estimated probability of a spurious zero $\tilde{b}_{ij}$ to 0.
  6. 6.
    $\tilde{z}_{ij,2}=0,\tilde{z}_{ij,1}\in\left(0,1\right),\tilde{z}_{ij,0}\in\left(0,1\right)$: In this case there are no reported double zeros so I set the estimated probability of a true zero $\tilde{p}_{ij}$ to 0. I then solve the system of equations:
    $\tilde{z}_{ij,1}$ $=\widetilde{Pr}\left\{\mathrm{one\ observed\ zero}|\mathrm{no\ spurious\ zeros,\ observed\ zeros}<2\right\}$
    $=\frac{2\tilde{b}_{ij}\left(1-\tilde{b}_{ij}\right)}{2\tilde{b}_{ij}\left(1-\tilde{b}_{ij}\right)+\left(1-\tilde{b}_{ij}\right)^{2}}=\frac{2\tilde{b}_{ij}}{1+\tilde{b}_{ij}}$
    $\tilde{z}_{ij,0}$ $=\widetilde{Pr}\left\{\mathrm{no\ observed\ zeros}|\mathrm{no\ spurious\ zeros,\ observed\ zeros}<2\right\}$
    $=\frac{\left(1-\tilde{b}_{ij}\right)^{2}}{2\tilde{b}_{ij}\left(1-\tilde{b}_{ij}\right)+\left(1-\tilde{b}_{ij}\right)^{2}}=\frac{1-\tilde{b}_{ij}}{1+\tilde{b}_{ij}},$
    and find
    $\tilde{b}_{ij}=\frac{\tilde{z}_{ij,1}}{2-\tilde{z}_{ij,1}}.$
G.3 Measurement Error Variances
We can combine the model equations to find:
$\log\tilde{F}_{ij,t}^{1}$ $=\beta_{t}\log\mathrm{dist}_{ij}+\alpha_{i,t}^{\mathrm{orig}}+\alpha_{j,t}^{\mathrm{dest}}+\eta_{ij,t}+\varepsilon_{ij,t}^{1},\quad\tilde{F}_{ij,t}^{1}>0$
$\log\tilde{F}_{ij,t}^{2}$ $=\beta_{t}\log\mathrm{dist}_{ij}+\alpha_{i,t}^{\mathrm{orig}}+\alpha_{j,t}^{\mathrm{dest}}+\eta_{ij,t}+\varepsilon_{ij,t}^{2},\quad\tilde{F}_{ij,t}^{2}>0,$
for i,j=1,,ni,j=1,...,n and t=1,,Tt=1,...,T. Subtracting these two equations yields
$\log\tilde{F}_{ij,t}^{1}-\log\tilde{F}_{ij,t}^{2}=\varepsilon_{ij,t}^{1}-\varepsilon_{ij,t}^{2}\sim\mathcal{N}\left(0,2\varsigma_{ij}^{2}\right),\quad\tilde{F}_{ij,t}^{1}>0,\tilde{F}_{ij,t}^{2}>0,$
for i,j=1,,ni,j=1,...,n and t=1,,Tt=1,...,T. This suggests the estimator
$\tilde{\varsigma}_{ij}^{2}=\frac{\mathbb{I}\left\{\sum_{t=1}^{T}\mathbb{I}\left\{\tilde{F}_{ij,t}^{1}>0,\tilde{F}_{ij,t}^{2}>0\right\}>0\right\}}{\sum_{t=1}^{T}\mathbb{I}\left\{\tilde{F}_{ij,t}^{1}>0,\tilde{F}_{ij,t}^{2}>0\right\}}\frac{1}{2}\sum_{t=1}^{T}\mathbb{I}\left\{\tilde{F}_{ij,t}^{1}>0,\tilde{F}_{ij,t}^{2}>0\right\}\cdot\left(\log\tilde{F}_{ij,t}^{1}-\log\tilde{F}_{ij,t}^{2}\right)^{2}$
for i,j=1,,ni,j=1,...,n. So note that county-pairs with no entries with two positive flows will have an estimated measurement error variance of 0. Note that the estimator is unbiased even with access to one period of mirror trade data (assuming both flows are non-negative). Obtaining estimators for the measurement error variances is what walters2024empirical calls the estimation step.
G.4 Prior Means
For the calibration of $\left(\left\{\beta_{t}\right\},\left\{\alpha_{i,t}^{\mathrm{exp}}\right\},\left\{\alpha_{j,t}^{\mathrm{imp}}\right\}\right)$, I use $\tilde{F}_{ij,t}=\tilde{F}_{ij,t}^{1}$. We then know that
$\log\tilde{F}_{ij,t}\sim\mathcal{N}\left(\beta_{t}\log\mathrm{dist}_{ij}+\alpha_{i,t}^{\mathrm{orig}}+\alpha_{j,t}^{\mathrm{dest}},s_{ij}^{2}+\varsigma_{ij}^{2}\right),\quad\mathrm{for}\ \tilde{F}_{ij,t}>0,$ (16)
for i,j=1,,ni,j=1,...,n and t=1,,Tt=1,...,T. [^18] Alternatively, one could use the average of the two measurements $0.5\left(\tilde{F}_{ij,t}^{1}+\tilde{F}_{ij,t}^{2}\right)$. In this case Equation (16) changes to $\log\tilde{F}_{ij,t}\sim\mathcal{N}\left(\beta_{t}\log\mathrm{dist}_{ij}+\alpha_{i,t}^{\mathrm{orig}}+\alpha_{j,t}^{\mathrm{dest}},s_{ij}^{2}+0.5\cdot\varsigma_{ij}^{2}\right)$. Using maximum likelihood estimation, it follows that the prior mean parameters can be estimated from the within-period regressions
$\log\tilde{F}_{ij,t}=\beta_{t}\log\mathrm{dist}_{ij}+\alpha_{i,t}^{\mathrm{orig}}+\alpha_{j,t}^{\mathrm{dest}}+\zeta_{ij,t},\quad\mathrm{for}\ \tilde{F}_{ij,t}>0,$ (17)
for t=1,,Tt=1,...,T, with $\zeta_{ij,t}$ an error term. The estimated prior means are
$\tilde{\mu}_{ij,t}$ $=\left(\tilde{\beta}_{t}\log\mathrm{dist}_{ij}+\tilde{\alpha}_{i,t}^{\mathrm{orig}}+\tilde{\alpha}_{j,t}^{\mathrm{dest}}\right)\cdot\mathbb{I}\left\{\tilde{F}_{ij,t}>0\right\}$
$\quad+\frac{\mathbb{I}\left\{\sum_{s=1}^{T}\mathbb{I}\left\{\tilde{F}_{ij,s}>0\right\}>0\right\}}{\sum_{s=1}^{T}\mathbb{I}\left\{\tilde{F}_{ij,s}>0\right\}}\cdot\sum_{s=1}^{T}\left\{\tilde{\beta}_{s}\log\mathrm{dist}_{ij}+\tilde{\alpha}_{i,s}^{\mathrm{orig}}+\tilde{\alpha}_{j,s}^{\mathrm{dest}}\right\},$
for i,j=1,,ni,j=1,...,n and t=1,,Tt=1,...,T. Note that for zero flows, the prior mean is imputed using an across-period average, and $\tilde{\mu}_{ij,t}$ is only zero if $\tilde{F}_{ij,t}$ is zero in all time periods for that country pair.
G.5 Prior Variances
From Equation (16) it follows that the posterior variances can be estimated by
$\tilde{s}_{ij}^{2}$ $=\max\left\{\widetilde{\operatorname{Var}}\left(\log\tilde{F}_{ij,t}-\tilde{\mu}_{ij,t}|\tilde{F}_{ij,t}>0\right)-\tilde{\varsigma}_{ij}^{2},0\right\},$
for i,j=1,,ni,j=1,...,n. Here, I again impute across periods for zero flows. Obtaining estimators for the prior means and variances is what walters2024empirical calls the deconvolution step.
G.6 Shrinking Variance Estimates
To leverage country information and the fact that importers and exporters can differ in their reliability, and reduce the variability for $\left\{\tilde{\varsigma}_{ij}^{2}\right\}$ and $\left\{\tilde{s}_{ij}^{2}\right\}$, I fit the models
$\tilde{\varsigma}_{ij}^{2}=e^{\kappa_{i}^{\varsigma,\mathrm{orig}}+\kappa_{j}^{\varsigma,\mathrm{dest}}+u_{ij}^{\varsigma}}\quad\mathrm{and}\quad\tilde{s}_{ij}^{2}=e^{\kappa_{i}^{s,\mathrm{orig}}+\kappa_{j}^{s,\mathrm{dest}}+u_{ij}^{s}},$ (18)
for i,j=1,,ni,j=1,...,n, with $\kappa_{i}^{\varsigma,\mathrm{orig}}$, $\kappa_{j}^{\varsigma,\mathrm{dest}}$, $\kappa_{i}^{s,\mathrm{orig}}$ and $\kappa_{j}^{s,\mathrm{dest}}$ country-origin and country-destination fixed effects and $u_{ij}^{\varsigma}$ and $u_{ij}^{s}$ error terms. Then, rather than using $\tilde{\varsigma}_{ij}^{2}$ and $\tilde{s}_{ij}^{2}$ I will use the fitted values $\mathring{\varsigma}_{ij}^{2}=e^{\tilde{\kappa}_{i}^{\varsigma,\mathrm{orig}}+\tilde{\kappa}_{j}^{\varsigma,\mathrm{dest}}}$ and $\mathring{s}_{ij}^{2}=e^{\tilde{\kappa}_{i}^{s,\mathrm{orig}}+\tilde{\kappa}_{j}^{s,\mathrm{dest}}}$.
G.7 Posterior Draws
It follows that the estimated posterior distribution for the true flow between location $i$ and $j$, $F_{ij,t}$, given its noisy version, $\tilde{F}_{ij,t}$ is given by
$F_{ij,t}|\tilde{F}_{ij,t},\tilde{\vartheta}\sim\begin{cases}\begin{array}[]{c}Q_{ij}\cdot\delta_{0}+\left(1-Q_{ij}\right)\cdot e^{\mathcal{N}\left(\tilde{\mu}_{ij,t},\mathring{s}_{ij}^{2}\right)}\\ \exp\left\{\mathcal{N}\left(\frac{\mathring{s}_{ij}^{2}}{\mathring{s}_{ij}^{2}+\mathring{\varsigma}_{ij}^{2}}\log\tilde{F}_{ij,t}+\frac{\mathring{\varsigma}_{ij}^{2}}{\mathring{s}_{ij}^{2}+\mathring{\varsigma}_{ij}^{2}}\tilde{\mu}_{ij,t},\left(\frac{1}{\mathring{s}_{ij}^{2}}+\frac{1}{\mathring{\varsigma}_{ij}^{2}}\right)^{-1}\right)\right\}\end{array}&\begin{array}[]{c}\tilde{F}_{ij}=0\\ \tilde{F}_{ij}>0\end{array},\end{cases}$ (19)
for i,j=1,,ni,j=1,...,n and t=1,,Tt=1,...,T, where $Q_{ij}\sim\mathrm{Bern}\left(\frac{\tilde{p}_{ij}}{\tilde{p}_{ij}+\tilde{b}_{ij}\left(1-\tilde{p}_{ij}\right)}\right)$.
G.8 Diagnostics
From Equation (16), one can verify how reasonable the normality assumption on the prior and measurement error model is by comparing the histogram of the normalized residuals
$\left\{\frac{\log\tilde{F}_{ij,t}-\tilde{\mu}_{ij,t}}{\sqrt{\mathring{s}_{ij}^{2}+\mathring{\varsigma}_{ij}^{2}}}\right\}_{i,j,t,\ \tilde{F}_{ij,t}>0}$
with the probability density function of a standard normal distribution. To further check the reasonableness of the gravity prior, we can look at the adjusted R-squared of the gravity regressions in Equation (17), and, following allen201813, plot the log flows against the log distance, after partitioning out the origin and destination fixed effects.
G.9 Computational Implementation Details
In the case where for all years one country reports only positive flows and the other country reports only NAs, I replace the NAs by the positive flows. After this initial replacement step, I replace the remaining NAs by zeros.
Appendix H Details for Armington Model
H.1 Derivation of System of Equations for $\left\{Y_{i}^{\mathrm{prop}}\right\}$
Rearranging Equation (10) and recalling that $\lambda_{ij}=F_{ij}/E_{j}$ yields:
$\lambda_{ij}=\frac{\left(\tau_{ij}Y_{i}\right)^{-\varepsilon}\chi_{ij}}{\sum_{k}\left(\tau_{kj}Y_{k}\right)^{-\varepsilon}\chi_{kj}},\qquad i,j=1,...,n.$ (20)
Next, plugging in Equations (11) and (20) into Equation (10) yields
$F_{ij}=\lambda_{ij}\left(1+\kappa_{j}\right)Y_{j},\qquad i,j=1,...,n.$
If we sum over $j$, we can use $Y_{i}=\sum_{\ell=1}^{n}F_{i\ell}$ to find
$Y_{i}=\sum_{j=1}^{n}\lambda_{ij}\left(1+\kappa_{j}\right)Y_{j},\qquad i=1,...,n.$ (21)
In the counterfactual equilibrium, Equation (21) should still hold. Because $\kappa_{i}$ is constant across equilibria for all $i$, this results in:
$Y_{i}^{\mathrm{cf,prop}}Y_{i}=\sum_{j=1}^{n}\lambda_{ij}^{\mathrm{cf,prop}}\lambda_{ij}\left(1+\kappa_{j}\right)Y_{j}^{\mathrm{cf,prop}}Y_{j},\qquad i=1,...,n.$ (22)
Similarly, Equation (20) should still hold in equilibrium. Using that $\chi_{ij}$ is constant across equilibria for all $i,j$, we find
$\lambda_{ij}^{\mathrm{cf,prop}}$ $=\frac{1}{\lambda_{ij}}\frac{\left(\tau_{ij}^{\mathrm{cf,prop}}\tau_{ij}Y_{i}^{\mathrm{cf,prop}}Y_{i}\right)^{-\varepsilon}\chi_{ij}}{\sum_{k}\left(\tau_{kj}^{\mathrm{cf,prop}}\tau_{kj}Y_{k}^{\mathrm{cf,prop}}Y_{k}\right)^{-\varepsilon}\chi_{kj}}$
$=\frac{1}{\lambda_{ij}}\frac{\left(\tau_{ij}^{\mathrm{cf,prop}}Y_{i}^{\mathrm{cf,prop}}\right)^{-\varepsilon}\frac{\left(\tau_{ij}Y_{i}^{\mathrm{}}\right)^{-\varepsilon}\chi_{ij}}{\sum_{\ell}\left(\tau_{\ell j}Y_{\ell}\right)^{-\varepsilon}\chi_{\ell j}}}{\sum_{k}\left(\tau_{kj}^{\mathrm{cf,prop}}Y_{k}^{\mathrm{cf,prop}}\right)^{-\varepsilon}\frac{\left(\tau_{kj}Y_{k}\right)^{-\varepsilon}\chi_{kj}}{\sum_{\ell}\left(\tau_{\ell j}Y_{\ell}\right)^{-\varepsilon}\chi_{\ell j}}}$
$=\frac{\left(\tau_{ij}^{\mathrm{cf,prop}}Y_{i}^{\mathrm{cf,prop}}\right)^{-\varepsilon}}{\sum_{k}\lambda_{kj}\left(\tau_{kj}^{\mathrm{cf,prop}}Y_{k}^{\mathrm{cf,prop}}\right)^{-\varepsilon}},\qquad i,j=1,...,n.$ (23)
Finally, combining Equations (22) and (23) yields the desired expression
$Y_{i}^{\mathrm{cf,prop}}Y_{i}=\sum_{j}\frac{\left(\tau_{ij}^{\mathrm{cf,prop}}Y_{i}^{\mathrm{cf,prop}}\right)^{-\varepsilon}}{\sum_{k}\lambda_{kj}\left(\tau_{kj}^{\mathrm{cf,prop}}Y_{k}^{\mathrm{cf,prop}}\right)^{-\varepsilon}}\lambda_{ij}\left(1+\kappa_{j}\right)Y_{j}^{\mathrm{cf,prop}}Y_{j},\qquad i=1,...,n.$
H.2 Results for Other Countries
Figure 9 reproduces Figure 1 for all 76 countries in the sample.
H.3 Calibration Procedure and Computational Details
The default approach from Section 4 can be applied. For the empirical Bayes step, the calibration of $\vartheta$, we can use the mirror trade data setting from Section 4.2.2. To construct $\left\{F_{ij}\right\}$, I use the mirror trade data for bilateral flows $\left\{F_{ij}\right\}_{i\neq j}$ and the trade flow data from waugh2010international for own-country flows $\left\{F_{ii}\right\}$. Because the mirror trade data report zero bilateral trade flows for Belgium, I exclude it from the analysis, resulting in a sample of 76 countries.
Appendix I Appendix Figures
[…]
Figure 5: Plot to compare the normalized residuals with the probability density function of a standardized normal distribution to check whether the normality assumption for the prior is reasonable for adao2017nonparametric.
Figure 6: Plot that follows allen201813 to check whether the gravity model is reasonable for log trade flows in 2011 from adao2017nonparametric.
Figure 7: Plot to compare the normalized residuals with the probability density function of a standardized normal distribution to check whether the normality assumption for the prior is reasonable for allen2022welfare.
Figure 8: Plot that follows allen201813 to check whether the gravity model is reasonable for log traffic flows from allen2022welfare.
[…]